### v2

# Dial-In LLM: Human-Aligned LLM-in-the-loop Intent Clustering for Customer Service Dialogues

Abstract
Discovering customer intentions in dialogue conversations is crucial for automated service agents. Yet, existing intent clustering methods often fail to align with human perceptions due to the heavy reliance on embedding distance metrics and sentence embeddings. To address these limitations, we propose integrating the semantic understanding capabilities of LLMs into an LLM-in-the-loop (LLM-ITL) intent clustering framework. Specifically, this paper (1) investigates the effectiveness of fine-tuned LLMs in semantic coherence evaluation and intent cluster naming, achieving over 95% accuracy; (2) designs an LLM-ITL clustering algorithm that facilitates the iterative discovery of coherent intent clusters; and (3) proposes task-specific techniques tailored for customer service dialogue intent clustering. Since existing English benchmarks pose limited semantic diversity and intent labels, we introduced a comprehensive Chinese dialogue intent dataset, comprising over 100,000 real customer service calls and 1,507 human-annotated intent clusters. The proposed approaches significantly outperformed LLM-guided baselines, achieving notable improvements in clustering quality and a 12% boost in the downstream intent classification task. Combined with several best practices, our findings highlight the potential of LLM-in-the-loop techniques for scalable and human-aligned problem-solving. Sample code and datasets are available at: anonymous.4open.science/r/Dial-in-LLM.
Dial-In LLM: Human-Aligned LLM-in-the-loop Intent Clustering for Customer Service Dialogues
Mengze Hong1  Di Jiang2thanks: Corresponding author: dijiang@webank.com  Yuanfeng Song2  Lu Wang3  Wailing Ng1 Yanjie Sun1  Chen Jason Zhang1  Qing Li1 1Hong Kong Polytechnic University  2AI Group, WeBank Co., Ltd  3Shenzhen University
1 Introduction
Intent discovery is an essential task in many NLP applications, such as designing AI-driven dialogue systems Hengst et al. (2024) and learning utterance patterns Ghosal et al. (2020). These systems often rely on an intent classification model that assigns intent labels. However, obtaining labeled data for training such supervised models is difficult, especially for dialogue data where topics evolve and contexts change Hengst et al. (2024); Hong et al. (2024). This motivates the need for robust intent discovery methods that can adapt to dynamic dialogue variations and applications.
Figure 1: Comparison of embedding-distance metric and human perception: cosine similarity failed to identify the same intentions under linguistic diversity (top), as well as distinguish different intentions under semantic diversity (bottom).
Intent clustering, which has evolved from unsupervised text clustering, offers significant advantages in automatically discovering thematic relationships Gung et al. (2023). Although text clustering is widely studied, existing research mainly concentrates on developing meaningful representations for text sentences Yin et al. (2021), treating it as a conventional clustering problem without addressing distinctive characteristics such as linguistic and semantic diversity (see Figure 1). Moreover, clustering algorithms heavily rely on quantitative optimization of embedding distance metrics with encoded representations, which is believed to hinder human-interpretable evaluation and validation.
With the growing research interest in integrating Large Language Models (LLMs) into the problem-solving pipeline Zhang et al. (2023); Viswanathan et al. (2024), known as the LLM-in-the-loop (LLM-ITL) paradigm Hong et al. (2025), this paper proposes an LLM-ITL solution specifically designed for dialogue intent clustering. This approach emphasizes the use of intermediate clustering results with human-aligned LLM utilities, providing practical benefits to industrial applications. The main contributions of this work include:
  1. 1.
    We introduce the largest Chinese dialogue intent clustering dataset, notable for its high semantic diversity and a significant inclusion of noisy, out-of-domain intents. This dataset is derived from over 100,000 real-world customer service call sessions, comprising 1,507 intent clusters annotated by human experts. It fills a crucial gap in evaluation benchmarks for complex intent clustering and classification.
  2. 2.
    We demonstrate the effectiveness of fine-tuned LLMs in assessing the semantic coherence of intent clusters and accurately providing intent labels across various sampling strategies. This highlights two effective and efficient utilities for designing LLM-in-the-loop solutions tailored for intent clustering.
  3. 3.
    We propose an LLM-in-the-loop clustering paradigm under both context-free and context-aware settings, which effectively utilizes intermediate results during the clustering process. This method surpasses state-of-the-art baselines and offers significant practical advantages in downstream applications.
2 Related Work
This study is grounded in two key areas of research: LLM-guided text clustering and intent clustering.
LLM-guided Text Clustering.
The discussion of text clustering with LLMs integration emerged as a trending topic since 2023. For instance, Zhang et al. (2023) utilized instruction-tuned LLMs like ChatGPT to better align clustering granularity with user preferences by refining the embedding space, while Viswanathan et al. (2024) demonstrated that LLMs can significantly enhance clustering quality by improving input features, providing constraints during clustering, and performing post-corrections. Feng et al. (2024) proposed a simple approach to refine edge points with LLM, resulting in minor performance improvement. Recently, Hong et al. (2025) proposed the concept of LLM-in-the-loop machine learning, categorizing LLM integration into data-, model-, and task-centric approaches. A “human-in-the-loop” solution can be replicated by LLM utilities Chen et al. (2024), enabling convenient and cost-efficient industrial deployment.
Intent Clustering.
Intent clustering is a fundamental step in intent induction Chandrakala et al. (2024) and the development of intent classifiers Gung et al. (2023). While various techniques have been proposed concerning semi-supervised Kumar et al. (2022) and deep learning approaches Lin et al. (2020); Zhang et al. (2021), recent advancements in LLM-based intent clustering systems have shown significant practical benefits Liang et al. (2024). Notably, the IDAS method proposed by De Raedt et al. (2023) is distinguished by its innovative use of LLMs and in-context learning to generate descriptive utterance labels, thereby enhancing clustering performance. However, the effective use of LLMs continues to pose challenges, such as managing input length and ensuring model robustness Song et al. (2023).
Highlights.
Previous studies have primary focused on a data-centric approach, prioritizing the enhancement of input text representations. In contrast, this paper emphasizes on the importance of intermediate results during the clustering process, designing tailored solutions for human-aligned intent clustering. Additionally, rather than relying on public benchmark datasets with limited complexity, this work tackles a more realistic setting by clustering a noisy dataset transcribed from real customer service calls, offering practical solutions in handling dialogue intent clustering.
3 Preliminaries
3.1 Coherence Evaluation
Inspired by human behaviors in perceiving texts and making comparisons based on semantic meaning rather than word-level similarity Peter W. Foltz and Landauer (1998), this paper proposes the semantic consistency within a cluster, known as semantic coherence, as a superior metric and optimization target. The evaluation is simplified by formulating it as a binary classification problem, leading toward intuitive evaluation and easing the training of the LLM evaluator. Good clusters consist of sentences focused on a specific topic, while Bad clusters contain inconsistent or ambiguous intentions (see Table 6 for examples). This task is particularly challenging for traditional machine learning classifiers due to a lack of semantic understanding Mimno et al. (2011), necessitating fine-tuned LLMs Gu et al. (2025) for robust evaluation.
3.2 Cluster Naming
Giving each cluster a concise and meaningful name is essential for many downstream applications Pattnaik et al. (2024). In this paper, we introduce a novel naming convention, "Action-Objective," which is particularly effective for capturing dialogue intents that are typically topic-oriented (e.g., insurance, loan) and involve distinct actions (e.g., inquire). Examples of human annotations are shown in Table 7, and a comparison of different naming conventions is presented in Table 8.
4 Proposed Methods
4.1 Context-Free Intent Clustering
Context-free intent clustering represents the simplest setting, which ignores task-related information such as dialogue roles and domain-specific knowledge. The performance of the context-free approach provides a general expectation of effectiveness when applying the proposed methods to different datasets and tasks.
4.1.1 LLM-in-the-loop Intent Clustering
With the introduction of cluster-level semantic coherence evaluation, individual clusters that are identified to be badly clustered can be further refined, enabling the iterative discovery of high-quality clusters. The proposed method is presented in Algorithm 1. To determine the appropriate number of clusters, which is often unknown in real-world applications, a candidate space of cluster numbers $\bm{N}$ is predefined. In each iteration, text embeddings $\bm{t}$ are input into a clustering algorithm $\bm{F}(\bm{t},n_{i})=\bm{c_{n_{i}}}$ for each candidate cluster number $n_{i}$. The coherence of the clustering results is then evaluated by a fine-tuned LLM, and the cluster assignment yielding the highest number of good clusters is selected as the intermediate result. The good clusters in this result are retained, and the sentences in the bad clusters are re-clustered in the subsequent iteration, ensuring that the final result consists mostly of high-quality clusters. The final clusters are labeled with LLM naming to improve the interpretability and practical values.
Determining the optimal number of clusters in a noisy text corpus is often challenging. Here, we propose a local search heuristic that maximizes the “good/bad” ratio at each iteration:
$\max_{n_{k}\in\bm{N}}\quad\frac{\#\,\bm{c_{n_{k}}}|\,g_{k}=good}{\#\,\bm{c_{n_{k}}}|g_{k}=bad}$
where $n_{k}$ denotes the best cluster number at the $k$th iteration. This enables the automatic discovery of suitable cluster numbers in a step-by-step manner. The search space $\bm{N}$ should chosen carefully to balance accuracy and efficiency; thus, we propose a search space pruning strategy, as illustrated in Section 6.4.2 for more details.
Algorithm 1 LLM-in-the-loop Intent Clustering
1:sentences, $LLM_{eval}$, $LLM_{name}$, cluster numbers $\bm{N}=[n_{1},\dots,n_{k}]$
2:set of clusters $\bm{C}$ with intent labels $\bm{L}$
3:$\bm{t}\leftarrow\text{SentenceTransformer}(sentences)$
4:$\bm{C}\leftarrow\emptyset$
5:for each $n_{i}\in N$ do
6:     initial clusters $\bm{c_{n_{i}}}\leftarrow\bm{F}(\bm{t},n_{i})$
7:     $\bm{g_{i}}\leftarrow[LLM_{eval}(c_{1}),\dots,LLM_{eval}(c_{n_{i}})]$
8:end for
9:$i\leftarrow\mathop{\arg\max}_{i}\frac{\#\,\bm{c_{n_{i}}}|\,g_{i}=good}{\#\,\bm{c_{n_{i}}}|g_{i}=bad}$
10:$\bm{C}\leftarrow\bm{C}\oplus\bm{c_{n_{i}}}\;|\,g_{i}=good$
11:if #$\#\,\bm{c_{n_{i}}}|\,g_{i}=bad>\epsilon$ then
12:     reverse to 3 with $\bm{t}\leftarrow\bm{c_{n_{i}}}\;|\,g_{i}=bad$
13:end if
[…]
4.1.2 Post-Correction with Intent Labels
From an initial exploration, multiple clusters representing similar intents may emerge in the final result (see Table 9). This occurs because, as the iterative process progresses, the number of poorly clustered sentences decreases, resulting in smaller and less representative clusters being formed. The challenge of embedding distance prevents the grouping of these clusters, motivating post-correction by merging similar clusters based on intent labels.
Unlike Viswanathan et al. (2024) and Feng et al. (2024), which directly query LLMs the correctness of cluster assignments (i.e., whether the cluster assignment of a data point is correct), we leverage the cluster naming capability of LLMs to label each cluster with its underlying intent. These labels $L(c_{i})$ can be merged when $\text{Dist}(L(c_{i}),L(c_{j}))<\theta$, where $\theta$ is a predefined merge threshold and Dist is a distance metric. As the intent label is a concrete and concise summary of the whole cluster, the embeddings of these labels are believed to retain semantic information and are suitable for merging similar clusters.
4.2 Context-Aware Intent Clustering
From a practical perspective, dialogue datasets are often accompanied by domain-specific features or use-case scenarios (e.g., customer service calls, group discussions). This motivates the incorporation of contextual information into the clustering process. However, the discovery of suitable knowledge remains a challenging task; thus, a knowledge discovery process is proposed based on LLM naming, where intermediate clustering assignments are utilized to achieve improved clustering results.
The customer service dialogues typically involve only the customer and the service agent Lin et al. (2022). The classification task that assigns a sentence to its associated role is relatively simple with labeled training data. To provide an unsupervised solution, the sentence roles are determined based on the labeled intent clusters. By heuristics, sentences within clusters labeled as “inquire-” or similar actions are mostly sent from the customer, and vice versa. A two-step approach is proposed: initially, intermediate clustering results are obtained using K-means or a context-free approach, and these results are divided based on intent labels, forming two clusters with distinct roles. In the second step, the sentences corresponding to each role are clustered again to achieve more refined outcomes, allowing for a clear separation of customer intent from service agent intent.
5 Experiment
5.1 Dataset
The proposed dataset contains 1,507 high-quality intent clusters manually annotated from over 100,000 customer service calls, comprising 55,085 distinct sentences with an average length of 17 Chinese characters per sentence (see Appendix A for annotation details) [^1] The full dataset will be open-sourced upon acceptance. Sample dataset can be found at GitHub.. Among these intents, 885 are identified as domain-specific queries (e.g., “inquire-insurance”), while 622 are categorized as out-of-domain (e.g., “provide-location”), representing general queries that might commonly arise in a customer service interaction. This dataset is the first Chinese benchmark for intent clustering and the largest of its kind compared to the existing English benchmark datasets shown in Table 1, which presents several new challenges such as lexical sparsity in short-text clustering, dependence on prior knowledge for context enrichment, and balancing accuracy with computational efficiency. The semantic diversity, calculated based on the average cosine distance between individual sentences and the cluster centroid Casanueva et al. (2022), depicts the complexity of this dataset.
Number of Number of Semantic
Dataset examples intents diversity
BANKING77 13,083 77 0.209
OOS 23,700 150 0.254
NLU++ 3,080 62 0.367
ours 55,085 1507 0.538
Table 1: Comparison of our proposed dataset with existing benchmark intent clustering datasets.
5.2 Metrics
Normalized Mutual Information (NMI) measures the degree of similarity between ground-truth and predicted clusters, scaled between 0 (no mutual information) and 1 (perfect correlation). However, researchers have claimed biased behavior in NMI Jerdee et al. (2024), necessitating further evaluation that reflects the true quality of the intent cluster from human perception.
Goodness score measures the proportion of good clusters evaluated by the fine-tuned LLM evaluator. For intermediate steps of iteration where the number of clusters is unknown, the good/bad ratio is used as an invariant measure of clustering quality:
$\text{goodness}_{i}=\frac{\text{\# good clusters}}{\text{\# bad clusters}}$
For the final clustering results, the percentage of good clusters among all clusters is reported. Note that the evaluation uses a different LLM evaluator than the one used for intermediate evaluation to ensure the fairness of the reported metric:
$\text{goodness}_{final}=\frac{\text{\# good clusters}}{\text{\# total clusters}}$
This metric offers a comprehensive understanding of the quality of the produced intent clusters, allowing for accurate cluster-level evaluation. Additionally, it can be easily deployed in applications without ground-truth annotations, thereby eliminating the need for human involvement.
Category Method NMI #Good
Base models K-Means 0.7899 94.8%
GMMs 0.7903 91.1%
Hierarchical* 0.8001 94.9%
Data-centric ClusterLLM Zhang et al. (2023) 0.7284 91.2%
IDAS De Raedt et al. (2023) 0.8109 93.6%
Keyphrase Viswanathan et al. (2024) 0.8371 94.5%
Context-Free LLM-ITL (random) 0.8202 97.7%
LLM-ITL (convex) 0.8208 97.8%
LLM-ITL + keyphrase 0.8378 96.4%
Context-Aware LLM-ITL + role 0.8679 97.2%
LLM-ITL + role + merge 0.8826 97.6%
Table 2: Performance of baselines and proposed methods in dialogue intent clustering. The best results are bolded.
6 Results and Discussions
6.1 LLM Utilities
{CJK}
UTF8gkai
LLM qwen7b qwen14b baichuan2-7 chatglm3-6b
Accuracy 96.25% 97.50% 89.17% 95.83%
Table 3: Performance of fine-tuned LLMs in evaluating semantic coherence of intent clusters.
Coherence Evaluation.
To enhance evaluation performance while preserving data privacy, four Chinese open-sourced LLMs Liu et al. (2024) are fine-tuned using LoRA Hu et al. (2022) on a training dataset comprising 1,772 intent clusters annotated as either “good” or “bad”. Table 3 presents performance of fine-tuned LLMs based on 480 unseen clusters. The results suggest that mainstream open-source LLMs can serve as robust evaluators for assessing the quality of intent clusters.
{CJK}
UTF8gkai
LLM qwen7b qwen14b baichuan2-7 chatglm3-6b
Accuracy 92.8% 94.3% 94.3% 94.4%
Table 4: Performance of fine-tuned LLMs in naming intent clusters.
Cluster Naming.
Similar to coherence evaluation, we manually annotated 2,500 clusters with corresponding intent labels to fine-tune and assess the utility of LLMs for cluster naming. Since labels are not unique, the accuracy is manually assessed by four human experts based on the alignment with the true labels in the dataset. The results in Table 4 suggest that fine-tuned LLMs have promising performance in providing intuitive names and aligning closely with human perceptions.
6.2 Intent Clustering Performance
Implementation.
The qwen14B model was utilized for both coherence evaluation and cluster naming. For all clustering methods, semantic-rich embeddings are obtained using the BAAI General Embeddings (BGE) model [^2] https://huggingface.co/BAAI/bge-large-zh-v1.5 Xiao et al. (2024). Among the clustering algorithms, hierarchical clustering demonstrated the highest NMI and goodness scores, thus serving as the base model for the proposed methods. Sampling techniques (i.e., random and convex) are implemented to select representative sentences from each cluster as input for LLM.
Among data-centric baselines, refining sentence embeddings through keyphrase expansion demonstrated the most improvement. However, deriving pairwise constraints for fine-tuning the embedding model, as proposed in ClusterLLM, proved ineffective due to the excessive number of clusters in the dataset. Interestingly, the evaluation of goodness does not always align with the NMI score, offering additional insights into cluster quality. For example, while keyphrase expansion improved the NMI score, the generated keyphrases could distort the meaning of the original sentence, resulting in less representative clusters and a slightly lower goodness score compared to the base model.
The proposed LLM-in-the-loop intent clustering, abbreviated as the “LLM-ITL” approach, demonstrated satisfactory improvement in NMI and significant enhancement in the goodness score. In a context-free setting, the integration of LLM evaluation and naming outperformed both ClusterLLM and IDAS baselines. By incorporating effective keyphrase expansion, the NMI improved further, highlighting the potential of integrating data-centric methods into the pipeline. Notably, the proposed method does not require prior knowledge of the number of clusters and conducts parameter searches during the clustering process, whereas the baseline methods preset the number of clusters at $n=1507$. This adaptability underscores its significant application value for extracting intent clusters from any noisy text corpora Akama et al. (2020).
Furthermore, context-aware approaches were evaluated with the goal of enhancing the task performance of customer service dialogue intent clustering. A consistent improvement was observed when incorporating dialogue role into the clustering process, resulting in a 4.48% increase in NMI without compromising the goodness score. This highlights the importance of task-centric design in effectively addressing industrial problems.
6.3 Application Performance
One objective of intent clustering is to create a high-quality labeled dataset for training intent classifiers that can effectively handle future data inputs Gung et al. (2023). To demonstrate the practical advantages of the proposed methods, we further evaluate their performance by training BERT classifiers (bert-base-chinese model) based on the clustered results. Compared to the training data produced by the baseline method, which achieved only 65% classification accuracy, the proposed LLM-in-the-loop method improved the accuracy by 12%, providing promising evidence of the improved cluster quality.
6.4 Ablation Studies
We present two ablation studies to derive best practices for industrial deployment using different sampling and hyperparameter tuning strategies.
6.4.1 Effective LLM Evaluation
Optimizing the input to LLMs is crucial for enhancing performance and reducing costs Song et al. (2025); Jiang et al. (2024). As shown in Figure 2, convex sampling outperforms random sampling in selecting representative sentences for coherence evaluation. Interestingly, as the sample time $t$ (for repeated evaluation) or the dimension of convex hull $d$ increases, performance consistently declines. This suggests that simpler, more efficient parameters lead to better accuracy. This pattern is verified through experiments with the ChatGLM model (see red line), which exhibits the same pattern.
To better assess the consistency and alignment between LLM judgments and human perceptions, we utilize the concept of crowdsourcing for collaborative coherence evaluation Zhang et al. (2025). The proposed LLM crowd consists of diverse entities Zhang et al. (2024) represented by various fine-tuned LLMs and uses a majority voting mechanism to aggregate their judgments Schoenegger et al. (2024). Based on the evaluation, we argue that a single fine-tuned LLM can effectively serve as a robust evaluator, as evidenced by the comparable performance to crowdsourced accuracy (cross markers in Figure 2) when using convex sampling.
6.4.2 Parameter Search Space Pruning
In practice, the search space for the number of clusters can be extensive, ranging from a few options to several hundred. To mitigate the computational costs associated with traditional model-based selection, we propose an LLM-native method for search space pruning, which predicts the optimal cluster number for subsequent iterations using logs from previous iterations (see Appendix 10). By pruning the search space and selecting the top 5 non-repetitive predicted solutions, we achieved an accuracy of 89.47% (see Table 5), indicating a high likelihood that the optimal cluster number is among these candidates. This approach substantially reduces redundant model fittings and enhances the efficiency of deploying the proposed methods.
{CJK}
UTF8gkai
Top1 Top2 Top3 Top4 Top5
Accuracy 26.32% 47.37% 73.68% 84.21% 89.47%
Table 5: Search space pruning by LLMs.
7 Conclusion
This paper tackles the dialogue intent clustering task by leveraging intermediate results through a human-aligned LLM-in-the-loop approach. Extensive experiments show that fine-tuned LLMs perform robustly in semantic coherence evaluations and cluster labeling. Both context-free and context-aware methods surpass existing LLM-guided clustering baselines, significantly improving cluster quality and downstream application performance. The ablation studies further provide practical insights for efficient industrial deployment. Future research is encouraged to enhance LLM reliability by integrating confidence estimation into the evaluation process, as well as employing advanced base algorithms and more sophisticated judgment formats to reveal additional insights.
Appendix A Data Collection and Annotation
This section details the three-stage data collection and annotation process for the proposed Chinese dialogue intent clustering dataset.
Data Collection.
The original dialogue dataset is sourced from audio transcriptions of customer service calls across several domains. Initially, the dataset contains over 120,000 calls. Strict filtering is applied to maintain data integrity and exclude confidential or sensitive content, resulting in 8,184 dialogues and 69,839 sentences. Subsequent data cleaning removes duplicate sentences, resulting in a final set of 55,085 unique sentences.
Data Annotation.
To create the intent clustering dataset, we recruited 15 human experts, each with over five years of experience in customer service call centers, to participate in the data annotation process:
  1. 1.
    K-means clustering was initially applied with $n=2000$, serving as a starting point for the data annotation.
  2. 2.
    Each expert assessed the initial clusters by inspecting the semantic coherence of the sentences. They were instructed: “Label the cluster as ’Good’ if all sentences share the same intent; otherwise, label it as ’Bad’. This resulted in 1283 good clusters and 717 bad clusters.
  3. 3.
    For the good clusters, annotators were asked to label the underlying intent using the naming convention “Action-Objective.” Five samples were provided as guides. After intent annotation, 1255 clusters with unique intentions were retained, and 28 clusters were merged due to replicated intent.
  4. 4.
    For the bad clusters, annotators reassigned each sentence to the appropriate intent cluster based on the annotated labels. Sentences that did not fit into any preexisting clusters were assigned to new clusters, and the same process was repeated as in Step 2.
Data Verification.
In the final stage, a separate group of 10 experts reviewed the annotated clusters for accuracy and consistency. Any discrepancies were resolved through consensus, ensuring the dataset’s reliability and validity for further analysis. The final dataset consists of 1,507 high-quality intent clusters.
Appendix B Prompt Template and More Results for LLM Utilities
The prompt for coherence evaluation and cluster naming are translated into English for demonstration. For each input, a few-shot demonstration with five input-output pairs is provided to ensure the consistency of the output format and enhance understanding of the task.
Coherence Evaluation -  Your are a helpful assistant for sentence clustering. Based on the relevancy and common points of the following sentences in a cluster, classify the cluster as: “Good” or “Bad”. Only provide the label without any additional content. Example: input:[sentences] output:[label] input:{[sentences]} output:
Cluster Naming -  Your are a helpful assistant for sentence clustering. Based on the relevancy and common points of the following sentences in a cluster, summarize the cluster with an “Action-Objective” label. Only provide the label without any additional content. Example: input:[sentences] output:[label] input:{[sentences]} output:
Considering the resource constraints of industrial deployment, we evaluated the coherence performance using the OpenAI APIs: GPT-3.5, GPT-4, and GPT-4o, achieving accuracies of 89.58%, 93.54%, and 94.17%, respectively. These findings indicate that a smaller, fine-tuned LLM, such as qwen7b, can outperform these advanced models and potentially reduce the costs associated with API requests [^3] The fine-tuned LLMs for coherence evaluation and cluster naming will be released upon acceptance.. To highlight the importance of fine-tuning, we also tested a vanilla Qwen2.5 model without fine-tuning, which resulted in a significantly lower accuracy of 75.63%.
Model Sampling Method 1 2 3 4 5
qwen14b convex - 96.25% 95.42% 95.42% 95.21%
random (n=10) 95.83% 94.38% 93.33% 92.92% 92.92%
random (n=20) 94.58% 95.00% 94.38% 94.17% 94.17%
chatglm3-6b convex - 92.08% 90.83% 91.04% 90.42%
random (n=10) 92.29% 88.75% 85.83% 84.58% 85.21%
random (n=20) 90.42% 89.58% 89.38% 89.58% 88.98%
Figure 2: Performance comparison of sampling methods and hyperparameters for effective LLM evaluation.
{CJK}
UTF8gkai
Cluster Coherence Original Sentences English Translation
Good "给企业固定资产买保险,大约能投大约得投保多少呢", "就是假如我有一百万的企业固定资金买保险大概投保多少钱啊", "请问一下我想咨>询一下企业财产保险", "您好我想了解一下这个企业财产保险", "就了解一下这个企业财产保险","你们是财险", "是哦这个属于财险了对吧", "企业财产保险的", "这些都属于财产险对吗", "财险人工那你这是", 就是如果我要为我这个私营企业买这个保险需要什么手续", "财产险", "呃企业财产保险是以什么为保" "If I buy insurance for a company’s fixed assets, how much insurance will it cost?", "That is, if I have a million corporate fixed assets, how much will it cost to buy insurance?", "Excuse me, I would like to consult> Ask about corporate property insurance", "Hello, I want to know about this corporate property insurance", "Just want to know about this corporate property insurance", "You are a property insurance company", "Yes, this belongs to property insurance "Right?", "Enterprise property insurance", "These all belong to property insurance, right?", "What about property insurance workers?", What do I need if I want to buy this insurance for my private enterprise? Procedure","Property Insurance", "Well, what does corporate property insurance cover?"
Bad "就是连连续,就是一直一直保", "但是它连不上是怎么回事啊?", "哦就是主要是直接给公司转账对吧?", "接吗", "接也是吗?", "我直>接去", "你是直接直接用那个" , "就是从哪接过来再接回去","还是需要从哪儿连线这个宽带", "你直接给我说这些啊", "直接把", "哦直接", "嗯那个礼品是直接就发放了", "是你直接给我回复对吗", "直接到那里去", "嗯,最好直飞。", "请帮我连接+", "把钱打过去的话,我是直接打到那个证券公司", "直接就是" "It’s continuous, it’s always guaranteed", "But what’s wrong with it not being able to connect?", "Oh, it’s mainly to transfer money directly to the company, right?", "Yes", "Yes too?" ?", "I’ll go directly", "Are you using that directly" , "Just connect it from where you are and then connect it back", "Or do you need to connect to the broadband from somewhere", "You Just tell me this directly", "Just give it directly", "Oh directly", "Well, the gift was given out directly", "You replied to me directly, right?","Go there directly" , "Well, it’s best to fly directly.", "Please help me connect +", "If I call the money, I will call the securities company directly", "Directly"
Table 6: Example of “Good” and “Bad” intent clusters in coherence evaluation.
{CJK}
UTF8gkai
Cluster Name Original Sentences English Translation
询问-优惠 (Inquire-Promotion) ’那有什么优惠券什么之类的吗?’, ’是是怎么形式是优惠吗?’, ’还是不是优惠活动,是那个直接给我打我卡里吗 ,还是是什么优惠券儿啊?’, ’就比如新用户他有什么优惠券儿之类的吧!’, ’嗯你你那还有什么优惠的活动吗? 就是比较合适就是合适的动。’ ’Are there any coupons or anything like that? ’, ’What is the form of the discount? ’, ’Still, it’s not a promotion. Is it the one that directly charges my card? , or is it some kind of coupon? ’, ’For example, what coupons does a new user have? ’, ’Well, do you have any other discounts? It is more appropriate and appropriate to move. ’
解答-金额 (Answer-Amount) ’一共是三十一块二毛’, ’就每个月一百三十八’, ’对,一个月也就是四百百四五百块钱嘛, 给您自己做个积累。’, ’十二月份的话是用了三十三块九毛二’, ’对一个月一百三十八’ ’The total is thirty-one and twenty cents’, ’That’s one hundred and thirty-eight cents per month’, ’Yes, that’s four hundred, four hundred and five hundred yuan per month. Make an accumulation for yourself. ’, ’In December, it cost thirty-three dollars and ninety-two cents’, ’That’s one hundred and thirty-eight dollars a month’
Table 7: Example of cluster naming with "Action-Objective" style.
Dataset Naming Convention Example (sentence) Example (label)
NLU scenario-intent Send an email to Alex and write thank you. email sendemail
NLU++ list of keywords How long does it usually take to get a new pin? ["how_long","pin", "arrival","new"]
OOS objective Please tell me why my card was declined yesterday. card_declined
ours action-objective Well, do you have any other discounts? inquire-promotion
Table 8: Comparison of cluster naming convention in intent clustering datasets.
{CJK}
UTF8gkai
Cluster Name Clusters with Similar Intention English Translation
询问-意外事故 (Inquire-Accident) "假如被车碰了或者是被楼上的砖砸了一下", "给别人儿撞的意外", "啊撞到别人然后就是", "就平时有时候开车嘛可能会遇到这个", "嗯哦这种情况,那要是就是我自己不小心撞到了那个某个地方然后", "就是把其他的东西撞到了呀什么的", "那人生意,不是就是,如果是不小心在马上被车撞了的话", "撞到人了吧", "然后不小在行驶当中被别人损害就是说拿石头砸的呀然后", "什么被车撞了之类的,是吗", "被被被撞了,还是被什么一些什么意外事故了", "我把别人的车撞了是吧" "If I were hit by a car or hit by a brick from upstairs," "Accidentally hit by someone else," "Ah, hit someone else and then," "Just sometimes when driving, you might encounter this," "Well, if I’m not careful and hit some place myself," "Just hit something else or something," "That’s a human affair, not just, if it’s accidentally hit by a car on the road," "Hit someone, right?" "Then not small in the process of driving, being damaged by someone else, say, hit with a stone, and then," "What, hit by a car or something, right?" "I hit someone else’s car, right?"
询问-意外死亡 (Inquire-Accident Death) "哦猝死,那猝死算意外吗", "啊,那我想知道那个猝死的话,算是意外死亡吗", "算意外死亡吗", "那如果猝死算是意外死亡吗?", "猝死也算意外事吧", "那那猝死是意外死亡吗如果是猝死的话是", "猝死算是意外死亡吗", "猝死算意外死亡吗", "嗯,那个猝死,猝死属于意外意外险吗,意外死亡吗", "那猝死的话,算意外死亡吗?", "那那那那个就是那个猝死算是意外死亡吗" "Oh sudden death, is sudden death considered accidental?" "Ah, I want to know if sudden death is considered accidental death?" "Is it considered accidental death?" "If sudden death is considered accidental death?" "Sudden death is also an accident, right?" "Is sudden death considered accidental death if it is sudden death?" "Is sudden death considered accidental death?" "Is sudden death considered accidental death?" "Well, that sudden death, does sudden death fall under accidental insurance, accidental death?" "Is sudden death considered accidental death?" "Is that, that, that sudden death considered accidental death?"
Table 9: Example of two high-quality intent clusters with similar intentions.
Epoch 1th 2th 3th
n_cluster good bad rate n_cluster good bad rate n_cluster good bad rate
20 20548 34537 0.595 20 0 2901 0.0 20 9 438 0.021
50 24100 30985 0.778 50 372 2529 0.147 50 70 377 0.186
100 32292 22793 1.417 100 804 2097 0.383 100 79 368 0.215
Best 1600 52184 2901 17.988 800 2454 447 5.490 200 86 361 0.238
Table 10: Example of log record for iterative intent clustering.

### v3

@@
 Discovering customer intentions in dialogue conversations is crucial for automated service agents.
-Yet, existing intent clustering methods often fail to align with human perceptions due to the heavy reliance on embedding distance metrics and sentence embeddings.
-To address these limitations, we propose integrating the semantic understanding capabilities of LLMs into an LLM-in-the-loop (LLM-ITL) intent clustering framework.
-Specifically, this paper (1) investigates the effectiveness of fine-tuned LLMs in semantic coherence evaluation and intent cluster naming, achieving over 95% accuracy; (2) designs an LLM-ITL clustering algorithm that facilitates the iterative discovery of coherent intent clusters; and (3) proposes task-specific techniques tailored for customer service dialogue intent clustering.
-Since existing English benchmarks pose limited semantic diversity and intent labels, we introduced a comprehensive Chinese dialogue intent dataset, comprising over 100,000 real customer service calls and 1,507 human-annotated intent clusters.
-The proposed approaches significantly outperformed LLM-guided baselines, achieving notable improvements in clustering quality and a 12% boost in the downstream intent classification task.
-Combined with several best practices, our findings highlight the potential of LLM-in-the-loop techniques for scalable and human-aligned problem-solving.
-Sample code and datasets are available at: anonymous.4open.science/r/Dial-in-LLM.
+However, existing intent clustering methods often fail to align with human perceptions due to a heavy reliance on embedding distance metrics and a tendency to overlook underlying semantic structures.
+This paper proposes an LLM-in-the-loop (LLM-ITL) intent clustering framework, integrating the semantic understanding capabilities of LLMs into conventional clustering algorithms.
+Specifically, this paper (1) investigates the effectiveness of fine-tuned LLMs in semantic coherence evaluation and intent cluster naming, achieving over 95% accuracy aligned with human judgments; (2) designs an LLM-ITL framework that facilitates the iterative discovery of coherent intent clusters and the optimal number of clusters; and (3) proposes context-aware techniques tailored for customer service dialogue.
+As existing English benchmarks offer limited semantic diversity and intent groups, we introduce a comprehensive Chinese dialogue intent dataset, comprising over 100k real customer service calls and 1,507 human-annotated intent clusters.
+The proposed approaches significantly outperform LLM-guided baselines, achieving notable enhancements in clustering quality and lower computational cost.
+Combined with several best practices, our findings highlight the potential of LLM-in-the-loop techniques for scalable and human-aligned intent clustering.
 Dial-In LLM: Human-Aligned LLM-in-the-loop Intent Clustering for Customer Service Dialogues
-Mengze Hong1  Di Jiang2thanks: Corresponding author: dijiang@webank.com  Yuanfeng Song2  Lu Wang3  Wailing Ng1 Yanjie Sun1  Chen Jason Zhang1  Qing Li1 1Hong Kong Polytechnic University  2AI Group, WeBank Co., Ltd  3Shenzhen University
+Mengze Hong1, Wailing Ng1, Chen Jason Zhang1 Yuanfeng Song2, Di Jiang2thanks: Corresponding author: dijiang@webank.com 1Hong Kong Polytechnic University 2AI Group, WeBank Co., Ltd
 1 Introduction
-Intent discovery is an essential task in many NLP applications, such as designing AI-driven dialogue systems Hengst et al.
-(2024) and learning utterance patterns Ghosal et al.
+Figure 1: Comparison of embedding-distance metric and human perception: cosine similarity failed to identify the same intentions under diverse expression (top), and to distinguish distinct intentions under similar expressions (bottom).
+Intent discovery is a crucial task in numerous NLP applications, including the design of AI-driven dialogue systems Hengst et al.
+(2024) and the learning of utterance patterns Ghosal et al.
 (2020).
-These systems often rely on an intent classification model that assigns intent labels.
-However, obtaining labeled data for training such supervised models is difficult, especially for dialogue data where topics evolve and contexts change Hengst et al.
-(2024); Hong et al.
+While intent clustering techniques are extensively studied to automatically identify thematic relationships within text corpora Gung et al.
+(2023), existing research primarily focuses on developing meaningful sentence representations and relies on embedding distance metrics for optimization Yin et al.
 (2021).
-This motivates the need for robust intent discovery methods that can adapt to dynamic dialogue variations and applications.
-Figure 1: Comparison of embedding-distance metric and human perception: cosine similarity failed to identify the same intentions under linguistic diversity (top), as well as distinguish different intentions under semantic diversity (bottom).
-Intent clustering, which has evolved from unsupervised text clustering, offers significant advantages in automatically discovering thematic relationships Gung et al.
+This approach often overlooks the distinctive characteristics of textual information, such as linguistic patterns and semantic diversity (see Figure 1), and restricts human-aligned evaluation and validation of clustering quality Vinh et al.
 (2009).
-Although text clustering is widely studied, existing research mainly concentrates on developing meaningful representations for text sentences Yin et al.
-(2021), treating it as a conventional clustering problem without addressing distinctive characteristics such as linguistic and semantic diversity (see Figure 1).
-Moreover, clustering algorithms heavily rely on quantitative optimization of embedding distance metrics with encoded representations, which is believed to hinder human-interpretable evaluation and validation.
-With the growing research interest in integrating Large Language Models (LLMs) into the problem-solving pipeline Zhang et al.
+This issue is particularly pronounced in languages with rich semantics, such as Chinese, where two seemingly similar sentences can convey entirely different meanings, and vice versa Chen (1993).
+With growing research interest in integrating Large Language Models (LLMs) into the problem-solving pipeline Hong et al.
+(2025), LLM-guided clustering techniques have emerged, demonstrating superior performance over traditional machine learning algorithms.
+While these methods effectively incorporate language understanding into the clustering process, they primarily focus on data preprocessing, querying LLMs for embedding refinement or data augmentation Zhang et al.
 (2023); Viswanathan et al.
-(2024), known as the LLM-in-the-loop (LLM-ITL) paradigm Hong et al.
-(2025), this paper proposes an LLM-ITL solution specifically designed for dialogue intent clustering.
-This approach emphasizes the use of intermediate clustering results with human-aligned LLM utilities, providing practical benefits to industrial applications.
-The main contributions of this work include:
+(2024).
+This approach represents a surface-level integration, potentially missing the benefits of LLMs in providing semantic-driven guidance to the clustering process and the quality of the final results.
+In this paper, we introduce an LLM-in-the-loop (LLM-ITL) intent clustering method, designed to facilitate the iterative discovery of coherent intent clusters.
+Our approach adeptly utilizes intermediate clustering results by incorporating human-aligned LLM utilities, offering computationally efficient and human-interpretable intent clustering.
+The key contributions of this work are:
   1. 1.
@@
   3. 3.
-    We propose an LLM-in-the-loop clustering paradigm under both context-free and context-aware settings, which effectively utilizes intermediate results during the clustering process.
-This method surpasses state-of-the-art baselines and offers significant practical advantages in downstream applications.
+    We propose an LLM-in-the-loop clustering algorithm to derive an optimal set of clusters in scenarios where the number of clusters is unknown.
+This approach effectively leverages intermediate results generated during the clustering process, outperforming state-of-the-art baselines and providing practical benefits for downstream applications.
+Additionally, the discussion on data sampling and LLM-based crowdsourcing further improves performance.
 2 Related Work
@@
 LLM-guided Text Clustering.
-The discussion of text clustering with LLMs integration emerged as a trending topic since 2023.
-For instance, Zhang et al.
-(2023) utilized instruction-tuned LLMs like ChatGPT to better align clustering granularity with user preferences by refining the embedding space, while Viswanathan et al.
-(2024) demonstrated that LLMs can significantly enhance clustering quality by improving input features, providing constraints during clustering, and performing post-corrections.
-Feng et al.
-(2024) proposed a simple approach to refine edge points with LLM, resulting in minor performance improvement.
+The integration of LLMs into text clustering has become increasingly prominent since 2023.
+Zhang et al.
+(2023) introduced ClusterLLM, an innovative approach that leverages instruction-tuned LLMs like ChatGPT to refine sentence embedding spaces through pairwise preference questions, aligning clustering granularity with user preferences.
+Viswanathan et al.
+(2024) highlighted the enhancement of clustering quality by LLMs through feature improvement, the imposition of constraints during clustering, and post-correction processes.
+Additionally, Feng et al.
+(2024) proposed refining edge points with LLMs, which led to notable performance gains.
 Recently, Hong et al.
-(2025) proposed the concept of LLM-in-the-loop machine learning, categorizing LLM integration into data-, model-, and task-centric approaches.
-A “human-in-the-loop” solution can be replicated by LLM utilities Chen et al.
-(2024), enabling convenient and cost-efficient industrial deployment.
+(2025) conceptualized LLM-in-the-loop machine learning, classifying LLM integration into data-, model-, and task-centric strategies.
+This method mirrors a “human-in-the-loop” approach Chen et al.
+(2024), facilitating effective LLM integration in machine learning applications.
 Intent Clustering.
@@
 (2023).
-While various techniques have been proposed concerning semi-supervised Kumar et al.
-(2022) and deep learning approaches Lin et al.
+While semi-supervised Kumar et al.
+(2022) and deep learning methods Lin et al.
 (2020); Zhang et al.
-(2021), recent advancements in LLM-based intent clustering systems have shown significant practical benefits Liang et al.
+(2021) have significantly advanced intent clustering, recent progress in LLM-based systems has highlighted their potential practical advantages Liang et al.
 (2024).
-Notably, the IDAS method proposed by De Raedt et al.
-(2023) is distinguished by its innovative use of LLMs and in-context learning to generate descriptive utterance labels, thereby enhancing clustering performance.
-However, the effective use of LLMs continues to pose challenges, such as managing input length and ensuring model robustness Song et al.
+The IDAS method effectively employs in-context learning to generate descriptive utterance labels and enhance sentence embeddings De Raedt et al.
+(2023).
+However, challenges such as computational cost and robustness remain unresolved Song et al.
 (2023).
 Highlights.
-Previous studies have primary focused on a data-centric approach, prioritizing the enhancement of input text representations.
-In contrast, this paper emphasizes on the importance of intermediate results during the clustering process, designing tailored solutions for human-aligned intent clustering.
-Additionally, rather than relying on public benchmark datasets with limited complexity, this work tackles a more realistic setting by clustering a noisy dataset transcribed from real customer service calls, offering practical solutions in handling dialogue intent clustering.
-3 Preliminaries
-3.1 Coherence Evaluation
+Previous studies have primarily focused on enhancing input text representations and the direct use of LLMs for cluster modification, imposing high computational cost and uncertainty.
+In contrast, this paper implements human-aligned LLM utilities to create a LLM-in-the-loop intent clustering process with emphasis on the intermediate results.
+Additionally, this work tackles a more realistic scenario by releasing a complex dataset transcribed from real customer service calls, offering practical insights and contributing to more meaningful evaluations in future intent clustering.
+Figure 2: Overview of proposed LLM-in-the-loop Dialogue Intent Clustering.
+3 Proposed Methods
+In this section, we first outline the design of human-aligned LLM utilities.
+Then, we introduce a comprehensive LLM-in-the-loop intent clustering framework (see Figure 2) and show how each LLM utility contributes to the clustering process.
+3.1 LLM Utilities
+Coherence Evaluation.
 Inspired by human behaviors in perceiving texts and making comparisons based on semantic meaning rather than word-level similarity Peter W.
 Foltz and Landauer (1998), this paper proposes the semantic consistency within a cluster, known as semantic coherence, as a superior metric and optimization target.
-The evaluation is simplified by formulating it as a binary classification problem, leading toward intuitive evaluation and easing the training of the LLM evaluator.
+The coherence evaluation is formulated as a binary classification problem, leading toward intuitive evaluation and easing the training of the LLM evaluator (see Section 5.6 for different formulations).
 Good clusters consist of sentences focused on a specific topic, while Bad clusters contain inconsistent or ambiguous intentions (see Table 9 for examples).
@@
 (2011), necessitating fine-tuned LLMs Gu et al.
-(2025) for robust evaluation.
-3.2 Cluster Naming
+(2025) for robust evaluation (see Section 5.1 for fine-tuning details).
+Cluster Naming.
 Giving each cluster a concise and meaningful name is essential for many downstream applications Pattnaik et al.
 (2024).
-In this paper, we introduce a novel naming convention, "Action-Objective," which is particularly effective for capturing dialogue intents that are typically topic-oriented (e.g., insurance, loan) and involve distinct actions (e.g., inquire).
-Examples of human annotations are shown in Table 7, and a comparison of different naming conventions is presented in Table 8.
-4 Proposed Methods
-4.1 Context-Free Intent Clustering
-Context-free intent clustering represents the simplest setting, which ignores task-related information such as dialogue roles and domain-specific knowledge.
-The performance of the context-free approach provides a general expectation of effectiveness when applying the proposed methods to different datasets and tasks.
-4.1.1 LLM-in-the-loop Intent Clustering
-With the introduction of cluster-level semantic coherence evaluation, individual clusters that are identified to be badly clustered can be further refined, enabling the iterative discovery of high-quality clusters.
-The proposed method is presented in Algorithm 1.
-To determine the appropriate number of clusters, which is often unknown in real-world applications, a candidate space of cluster numbers $\bm{N}$ is predefined.
-In each iteration, text embeddings $\bm{t}$ are input into a clustering algorithm $\bm{F}(\bm{t},n_{i})=\bm{c_{n_{i}}}$ for each candidate cluster number $n_{i}$.
-The coherence of the clustering results is then evaluated by a fine-tuned LLM, and the cluster assignment yielding the highest number of good clusters is selected as the intermediate result.
-The good clusters in this result are retained, and the sentences in the bad clusters are re-clustered in the subsequent iteration, ensuring that the final result consists mostly of high-quality clusters.
-The final clusters are labeled with LLM naming to improve the interpretability and practical values.
-Determining the optimal number of clusters in a noisy text corpus is often challenging. Here, we propose a local search heuristic that maximizes the “good/bad” ratio at each iteration:
-$\max_{n_{k}\in\bm{N}}\quad\frac{\#\,\bm{c_{n_{k}}}|\,g_{k}=good}{\#\,\bm{c_{n_{k}}}|g_{k}=bad}$
-where $n_{k}$ denotes the best cluster number at the $k$th iteration.
+In this paper, we introduce a novel naming convention, "Action-Objective," which is particularly effective for capturing dialogue intents that are typically topic-oriented (e.g., insurance, loan) and involve distinct actions (e.g., inquire, confirm).
+Examples of human annotations are shown in Table 10, and a comparison of different naming conventions is presented in Table 11, demonstrating the superiority of the proposed convention in capturing customer intention with clear goal-oriented expression.
+3.2 LLM-in-the-loop Iterative Intent Clustering with Coherence Evaluation
+Algorithm 1 LLM-in-the-loop Intent Clustering
+1:Unlabeled sentence corpus $\mathcal{S}$, embedding function $f_{\text{emb}}$, coherence evaluator $\mathcal{M}_{\text{eval}}$, candidate cluster numbers $N=\{n_{1},\dots,n_{k}\}$, threshold $\epsilon>0$, max iterations $T_{\text{max}}$
+2:Set of clusters $\mathcal{C}$
+3:$\mathcal{E}\leftarrow\{f_{\text{emb}}(s)\mid s\in\mathcal{S}\}$
+4:$\mathcal{S}^{(0)}\leftarrow\mathcal{S}$, $\mathcal{C}\leftarrow\emptyset$, $t\leftarrow 0$
+5:while $\frac{|\mathcal{S}^{(t)}|}{|\mathcal{S}|}>\epsilon$ and $t<T_{\text{max}}$ do
+6:     $\mathcal{E}^{(t)}\leftarrow\{f_{\text{emb}}(s)\mid s\in\mathcal{S}^{(t)}\}$
+7:     for each $n_{i}\in N$ do
+8:         $\mathcal{C}_{n_{i}}^{(t)}\leftarrow F(\mathcal{E}^{(t)},n_{i})$
+9:         $\mathbf{g}_{n_{i}}^{(t)}\leftarrow[\mathcal{M}_{\text{eval}}(C_{1}^{(t)}),\dots,\mathcal{M}_{\text{eval}}(C_{n_{i}}^{(t)})]$
+10:     end for
+11:     $n^{*}\leftarrow\arg\max_{n_{i}\in N}\frac{\sum_{j=1}^{n_{i}}\mathbb{I}\left(\mathbf{g}_{n_{i}}^{(t)}[j]=1\right)}{\sum_{j=1}^{n_{i}}\mathbb{I}\left(\mathbf{g}_{n_{i}}^{(t)}[j]=0\right)+1}$
+12:     $\mathcal{C}_{\text{good}}^{(t)}\leftarrow\{C_{j}\in\mathcal{C}_{n^{*}}^{(t)}\mid\mathbf{g}_{n^{*}}^{(t)}[j]=1\}$
+13:     $\mathcal{C}\leftarrow\mathcal{C}\cup\mathcal{C}_{\text{good}}^{(t)}$
+14:     $\mathcal{S}^{(t+1)}\leftarrow\mathcal{S}^{(t)}\setminus\bigcup_{C\in\mathcal{C}_{\text{good}}^{(t)}}C$
+15:     $t\leftarrow t+1$
+16:end while
+17:return $(\mathcal{C})$
+At iteration $t$, the current set of unassigned sentences is denoted as $\mathcal{S}^{(t)}$.
+A special case in the first iteration, where $\mathcal{S}^{(0)}=\mathcal{S}$ represents the entire set of unique sentences derived from the dialogue corpus.
+For each candidate cluster number $n_{i}\in N$, we compute:
+$\mathcal{C}_{n_{i}}^{(t)}=F(\mathcal{E}^{(t)},n_{i}),$
+where $F$ is a clustering function (e.g., K-means clustering), and $\mathcal{E}^{(t)}=\{\mathbf{e}_{s}\mid s\in\mathcal{S}^{(t)}\}$ representing the sentence embeddings.
+This results in $N$ distinct cluster assignments.
+Then, the semantic coherence of each cluster is evaluated using a fine-tuned LLM $\mathcal{M}_{\text{eval}}$:
+$\mathbf{g}_{n_{i}}^{(t)}=\left[\mathcal{M}_{\text{eval}}(C_{1}^{(t)}),\ldots,\mathcal{M}_{\text{eval}}(C_{n_{i}}^{(t)})\right],$
+where $\mathcal{M}_{\text{eval}}(C_{k})=1$ if the cluster is coherent (i.e., “good” cluster), else 0.
+While the number of clusters is often known with benchmark evaluation, it is challenging but necessary to implement mechanisms for determining the optimal number of clusters in a noisy text corpus.
+Here, we propose a local search heuristic that maximizes the “good/bad” ratio at each iteration.
+The optimal $n_{t}^{*}$ is selected as:
+$n_{t}^{*}=\arg\max_{n_{i}\in N}\frac{\sum_{j=1}^{n_{i}}\mathbb{I}\left(\mathbf{g}_{n_{i}}^{(t)}[j]=1\right)}{\sum_{j=1}^{n_{i}}\mathbb{I}\left(\mathbf{g}_{n_{i}}^{(t)}[j]=0\right)+1},$
+where $n_{t}$ denotes the best cluster number at iteration $t$.
 This enables the automatic discovery of suitable cluster numbers in a step-by-step manner.
-The search space $\bm{N}$ should chosen carefully to balance accuracy and efficiency; thus, we propose a search space pruning strategy, as illustrated in Section 6.4.2 for more details.
-Algorithm 1 LLM-in-the-loop Intent Clustering
-1:sentences, $LLM_{eval}$, $LLM_{name}$, cluster numbers $\bm{N}=[n_{1},\dots,n_{k}]$
-2:set of clusters $\bm{C}$ with intent labels $\bm{L}$
-3:$\bm{t}\leftarrow\text{SentenceTransformer}(sentences)$
-4:$\bm{C}\leftarrow\emptyset$
-5:for each $n_{i}\in N$ do
-6:     initial clusters $\bm{c_{n_{i}}}\leftarrow\bm{F}(\bm{t},n_{i})$
-7:     $\bm{g_{i}}\leftarrow[LLM_{eval}(c_{1}),\dots,LLM_{eval}(c_{n_{i}})]$
-8:end for
-9:$i\leftarrow\mathop{\arg\max}_{i}\frac{\#\,\bm{c_{n_{i}}}|\,g_{i}=good}{\#\,\bm{c_{n_{i}}}|g_{i}=bad}$
-10:$\bm{C}\leftarrow\bm{C}\oplus\bm{c_{n_{i}}}\;|\,g_{i}=good$
-11:if #$\#\,\bm{c_{n_{i}}}|\,g_{i}=bad>\epsilon$ then
-12:     reverse to 3 with $\bm{t}\leftarrow\bm{c_{n_{i}}}\;|\,g_{i}=bad$
-13:end if
-[…]
-4.1.2 Post-Correction with Intent Labels
-From an initial exploration, multiple clusters representing similar intents may emerge in the final result (see Table 9).
-This occurs because, as the iterative process progresses, the number of poorly clustered sentences decreases, resulting in smaller and less representative clusters being formed.
-The challenge of embedding distance prevents the grouping of these clusters, motivating post-correction by merging similar clusters based on intent labels.
-Unlike Viswanathan et al.
-(2024) and Feng et al.
-(2024), which directly query LLMs the correctness of cluster assignments (i.e., whether the cluster assignment of a data point is correct), we leverage the cluster naming capability of LLMs to label each cluster with its underlying intent.
-These labels $L(c_{i})$ can be merged when $\text{Dist}(L(c_{i}),L(c_{j}))<\theta$, where $\theta$ is a predefined merge threshold and Dist is a distance metric.
-As the intent label is a concrete and concise summary of the whole cluster, the embeddings of these labels are believed to retain semantic information and are suitable for merging similar clusters.
-4.2 Context-Aware Intent Clustering
+The search space $\bm{N}$ should be chosen carefully to balance accuracy and efficiency; thus, we propose a search space pruning strategy, as illustrated in Section 5.6 for more details.
+Finally, the “good” clusters in the optimal $\mathcal{C}_{n^{*}}^{(t)}$ are retained, and the remaining sentences will be refined in the next (i.e., $t+1$) iteration, enabling the iterative discovery of high-quality clusters.
+The proposed method is summarized in Algorithm 1, and the LLM integration is highlighted.
+3.3 Post-Correction with LLM-Generated Intent Labels
+Preliminary results in Table 12 suggest that multiple clusters may capture similar intents due to diminishing size of unassigned sentences $\mathcal{S}^{(t)}$ in later iterations, leading to smaller and less representative clusters.
+The embedding distances between sentences within clusters limit the natural consolidation of clusters with similar intents but different expressions Khan et al.
+(2020), necessitating a post-correction step to merge semantically aligned clusters.
+In contrast to previous methods that rely on direct LLM queries for validating individual cluster assignments Viswanathan et al.
+(2024); Feng et al.
+(2024), which are computationally expensive and inconsistent, we propose a context-aware approach, leveraging LLMs’ naming utility to robustly merge clusters based on intent labels.
+At the end of the iterative clustering process, each cluster $C_{k}$ receives an intent label from the fine-tuned LLM naming utility:
+$l_{k}=\mathcal{M}_{\text{name}}(C_{k}).$
+These labels concisely summarize the semantic information of each cluster and are mapped to the semantic space of sentence embeddings using the embedding function $f_{\text{emb}}$:
+$\mathbf{l}_{k}=f_{\text{emb}}(l_{k})\in\mathbb{R}^{d},\quad\|\mathbf{l}_{k}\|_{2}=1.$
+This normalization positions label embeddings $\mathbf{l}_{k}$ on the unit hypersphere $\mathbb{S}^{d-1}=\{\mathbf{x}\in\mathbb{R}^{d}\mid\|\mathbf{x}\|_{2}=1\}$, enabling accurate measurement of semantic relationships along the sphere’s surface rather than through straight-line distances Fletcher et al.
+(2004), thus facilitating precise intent similarity comparisons in high-dimensional space.
+Then, the clusters are structured into a semantic affinity graph $G=(V,E)$ to model relationships based on label similarity.
+Vertices $V=\{C_{1},C_{2},\ldots,C_{K}\}$ represent clusters, and the edges $E\subseteq V\times V$ are determined using hyperspherical geometry by computing the geodesic distance, which captures angular separation between label embeddings Fletcher et al.
+(2004):
+$\text{Dist}(\mathbf{l}_{i},\mathbf{l}_{j})=\arccos(\langle\mathbf{l}_{i},\mathbf{l}_{j}\rangle),$
+forming edges if $\text{Dist}(\mathbf{l}_{i},\mathbf{l}_{j})<\theta$, with $\theta$ is a predefined threshold ($\theta=0.8$).
+For each derived edge, a probabilistic criterion is designed to enhance the robustness of merging decisions, naturally modeling label embeddings as samples from a mixture of von Mises-Fisher distributions Banerjee et al.
+(2005):
+$p(\mathbf{l}_{k}\mid\boldsymbol{\mu}_{m},\kappa_{m})=Z_{d}(\kappa_{m})\exp(\kappa_{m}\langle\mathbf{l}_{k},\boldsymbol{\mu}_{m}\rangle),$
+where $\boldsymbol{\mu}_{m}=\mathbf{l}_{m}\in\mathbb{S}^{d-1}$ is the mean direction of the $m$-th intent embedding, $\kappa_{m}>0$ controls the distribution’s tightness, and the normalization constant $Z_{d}(\kappa_{m})$ ensures the density integrates to 1 over the hypersphere:
+$Z_{d}(\kappa)=\frac{\kappa^{d/2-1}}{(2\pi)^{d/2}I_{d/2-1}(\kappa)}.$
+Thus, the probability that clusters $C_{i}$ and $C_{j}$ share the same intent is computed for each edge as:
+$P(\text{same}\mid\mathbf{l}_{i},\mathbf{l}_{j})=\sum_{m=1}^{K}\pi_{m}p(\mathbf{l}_{i}\mid\boldsymbol{\mu}_{m},\kappa_{m})p(\mathbf{l}_{j}\mid\boldsymbol{\mu}_{m},\kappa_{m}),$
+with $\pi_{m}=1/K$ as uniform mixture weights.
+An edge is retained only if this probability exceeds a predefined threshold $\tau$ ($\tau=0.7$), minimizing inappropriate merges by accounting for uncertainty.
+Finally, clusters linked in the affinity graph are consolidated into connected components, forming a refined cluster assignment $\mathcal{C}^{\prime}=\{C^{\prime}_{1},C^{\prime}_{2},\ldots,C^{\prime}_{K^{\prime}}\}$ that eliminates redundancy, enhances interpretability, and maintains robust semantic alignment.
+Each merged cluster $C^{\prime}_{k}$ receives a new intent label reflecting its semantic content.
+3.4 Context-Aware Role Separation with LLM-Generated Intent Labels
 From a practical perspective, dialogue datasets are often accompanied by domain-specific features or use-case scenarios (e.g., customer service calls, group discussions).
 This motivates the incorporation of contextual information into the clustering process.
-However, the discovery of suitable knowledge remains a challenging task; thus, a knowledge discovery process is proposed based on LLM naming, where intermediate clustering assignments are utilized to achieve improved clustering results.
 The customer service dialogues typically involve only the customer and the service agent Lin et al.
@@
 The classification task that assigns a sentence to its associated role is relatively simple with labeled training data.
-To provide an unsupervised solution, the sentence roles are determined based on the labeled intent clusters.
+To provide an unsupervised solution, the sentence roles are naturally determined based on the LLM- intent label.
 By heuristics, sentences within clusters labeled as “inquire-” or similar actions are mostly sent from the customer, and vice versa.
-A two-step approach is proposed: initially, intermediate clustering results are obtained using K-means or a context-free approach, and these results are divided based on intent labels, forming two clusters with distinct roles.
-In the second step, the sentences corresponding to each role are clustered again to achieve more refined outcomes, allowing for a clear separation of customer intent from service agent intent.
+A two-step approach is proposed: initially, intermediate clustering results are obtained using previous approach, denoted as $C_{\text{inter}}=\{C_{1},C_{2},\ldots,C_{K}\}$, and these results are divided based on intent labels $l_{k}$, forming two groups with distinct roles, $R_{\text{customer}}$ and $R_{\text{agent}}$, where:
+$\begin{array}[]{c}R_{\text{customer}}=\{s\in S\mid l_{k}(s)\in\{\text{"inquire-"},\dots\}\},\\ R_{\text{agent}}=\{s\in S\mid l_{k}(s)\notin\{\text{"answer-"},\dots\}\}.\end{array}$
+In the second step, the sentences corresponding to each role are clustered again, denoted as $C^{\prime}_{\text{customer}}=\text{Cluster}(R_{\text{customer}})$ and $C^{\prime}_{\text{agent}}=\text{Cluster}(R_{\text{agent}})$, to achieve more refined outcomes, allowing for a clear separation of customer intent from service agent intent.
+Finally, the two groups of clusters are merged together, serving as the final or intermediate result to be further processed.
 4 Experiment
 4.1 Dataset
-The proposed dataset contains 1,507 high-quality intent clusters manually annotated from over 100,000 customer service calls, comprising 55,085 distinct sentences with an average length of 17 Chinese characters per sentence (see Appendix A for annotation details) [^1] The full dataset will be open-sourced upon acceptance.
+The proposed dataset contains 1,507 high-quality intent clusters manually annotated from over 100,000 realistic customer service calls, comprising 55,085 distinct sentences with an average length of 17 Chinese characters per sentence (see Appendix A for annotation details) [^1] The full dataset and fine-tuned models will be released upon acceptance.
 Sample dataset can be found at GitHub..
-Among these intents, 885 are identified as domain-specific queries (e.g., “inquire-insurance”), while 622 are categorized as out-of-domain (e.g., “provide-location”), representing general queries that might commonly arise in a customer service interaction.
-This dataset is the first Chinese benchmark for intent clustering and the largest of its kind compared to the existing English benchmark datasets shown in Table 1, which presents several new challenges such as lexical sparsity in short-text clustering, dependence on prior knowledge for context enrichment, and balancing accuracy with computational efficiency.
+Among these intents, 885 are identified as domain-specific queries (e.g., inquire-insurance), while 622 are categorized as out-of-domain (e.g., provide-location), representing general queries that commonly arise in a customer service interaction.
+This dataset is the first Chinese benchmark for intent clustering and the largest of its kind compared to the existing English benchmark datasets shown in Table 1.
+It presents several new challenges, including the lexical sparsity in short-text sentences, the dependence on contextual knowledge, and the difficulties in balancing accuracy with computational efficiency due to the excessively large intent size.
 The semantic diversity, calculated based on the average cosine distance between individual sentences and the cluster centroid Casanueva et al.
-(2022), depicts the complexity of this dataset.
+(2022), depicts the complexity of this dataset and motivates the development of semantic-guided techniques.
 Number of Number of Semantic
-Dataset examples intents diversity
-BANKING77 13,083 77 0.209
+Dataset sentences intents diversity
+BANK77 3080 77 0.209
 OOS 23,700 150 0.254
 NLU++ 3,080 62 0.367
+CLINC(I) 4,500 150 0.275
+MTOP(I) 4,386 102 0.234
+Massive(I) 2,974 59 0.351
 ours 55,085 1507 0.538
-Table 1: Comparison of our proposed dataset with existing benchmark intent clustering datasets.
+Table 1: Comparison of proposed dataset with existing intent clustering datasets.
+Category Method NMI #Good
+Baselines K-Means 0.7899 94.8%
+GMMs 0.7903 91.1%
+Hierarchical* 0.8001 94.9%
+LLM-Guided Clustering Baselines ClusterLLM Zhang et al. (2023) 0.7284 91.2%
+IDAS De Raedt et al. (2023) 0.8109 93.6%
+Keyphrase Viswanathan et al. (2024) 0.8371 94.5%
+LLMEdgeRefine Feng et al. (2024) 0.7411 87.2%
+Proposed Method (context-free) LLM-ITL (random) 0.8202 97.7%
+LLM-ITL (convex) 0.8208 97.8%
+LLM-ITL + keyphrase 0.8378 96.4%
+Proposed Method (context-aware) LLM-ITL + merge 0.8420 97.8%
+LLM-ITL + role 0.8679 97.2%
+LLM-ITL + role + merge 0.8826 97.6%
+Table 2: Performance of baselines and proposed methods in dialogue intent clustering. The best results are bolded.
+{CJK}
+UTF8gkai
+LLM qwen7b qwen14b baichuan2-7b chatglm3-6b
+Accuracy 96.25% 97.50% 89.17% 95.83%
+Table 3: Performance of fine-tuned LLMs in evaluating semantic coherence of intent clusters.
+{CJK}
+UTF8gkai
+LLM qwen7b qwen14b baichuan2-7b chatglm3-6b
+Accuracy 92.8% 94.3% 94.3% 94.4%
+Table 4: Performance of fine-tuned LLMs in naming intent clusters.
 4.2 Metrics
@@
 Additionally, it can be easily deployed in applications without ground-truth annotations, thereby eliminating the need for human involvement.
-Category Method NMI #Good
-Base models K-Means 0.7899 94.8%
-GMMs 0.7903 91.1%
-Hierarchical* 0.8001 94.9%
-Data-centric ClusterLLM Zhang et al. (2023) 0.7284 91.2%
-IDAS De Raedt et al. (2023) 0.8109 93.6%
-Keyphrase Viswanathan et al. (2024) 0.8371 94.5%
-Context-Free LLM-ITL (random) 0.8202 97.7%
-LLM-ITL (convex) 0.8208 97.8%
-LLM-ITL + keyphrase 0.8378 96.4%
-Context-Aware LLM-ITL + role 0.8679 97.2%
-LLM-ITL + role + merge 0.8826 97.6%
-Table 2: Performance of baselines and proposed methods in dialogue intent clustering. The best results are bolded.
+4.3 Implementations
+Four open-sourced Chinese LLMs [^2] The models are available at: Qwen2.5-7B; Qwen2.5-14B; Baichuan2-7B ; ChatGLM3-6b are fine-tuned using LoRA Hu et al.
+(2022) on 4 $\times$ Nvidia A100 GPUs.
+For coherence evaluation, a human-annotated training dataset of 1,772 intent clusters labeled as “good” or “bad” is used.
+For cluster naming, a training dataset of 2,500 clusters, each containing 20 sentences, is annotated with intent labels.
+Semantic-rich embeddings are generated using the BAAI General Embeddings (BGE) model [^3] https://huggingface.co/BAAI/bge-large-zh-v1.5 Xiao et al.
+(2024).
+The best-performing clustering algorithm, hierarchical clustering, and LLM utilities, qwen14b and chatglm3-6b, are selected for LLM-in-the-loop clustering.
+To reduce computational load, both random and convex sampling methods are implemented to select 20 representative sentences per cluster as input for the LLM.
 5 Results and Discussions
 5.1 LLM Utilities
-{CJK}
-UTF8gkai
-LLM qwen7b qwen14b baichuan2-7 chatglm3-6b
-Accuracy 96.25% 97.50% 89.17% 95.83%
-Table 3: Performance of fine-tuned LLMs in evaluating semantic coherence of intent clusters.
-Coherence Evaluation.
-To enhance evaluation performance while preserving data privacy, four Chinese open-sourced LLMs Liu et al.
-(2024) are fine-tuned using LoRA Hu et al.
-(2022) on a training dataset comprising 1,772 intent clusters annotated as either “good” or “bad”.
-Table 3 presents performance of fine-tuned LLMs based on 480 unseen clusters.
-The results suggest that mainstream open-source LLMs can serve as robust evaluators for assessing the quality of intent clusters.
-{CJK}
-UTF8gkai
-LLM qwen7b qwen14b baichuan2-7 chatglm3-6b
-Accuracy 92.8% 94.3% 94.3% 94.4%
-Table 4: Performance of fine-tuned LLMs in naming intent clusters.
-Cluster Naming.
-Similar to coherence evaluation, we manually annotated 2,500 clusters with corresponding intent labels to fine-tune and assess the utility of LLMs for cluster naming.
-Since labels are not unique, the accuracy is manually assessed by four human experts based on the alignment with the true labels in the dataset.
-The results in Table 4 suggest that fine-tuned LLMs have promising performance in providing intuitive names and aligning closely with human perceptions.
-6.2 Intent Clustering Performance
-Implementation.
-The qwen14B model was utilized for both coherence evaluation and cluster naming.
-For all clustering methods, semantic-rich embeddings are obtained using the BAAI General Embeddings (BGE) model [^2] https://huggingface.co/BAAI/bge-large-zh-v1.5 Xiao et al.
+Table 3 presents the performance of the fine-tuned LLM coherence evaluator tested on 480 unseen clusters.
+The results indicate that mainstream open-source LLMs can effectively serve as robust evaluators for assessing the quality of intent clusters and providing human-aligned judgments.
+For cluster naming, since labels are not unique, accuracy is manually evaluated by four human experts based on alignment with the true labels in the dataset.
+The results in Table 4 demonstrate that the fine-tuned LLMs show promising performance in generating intuitive names and adhering strictly to the predefined “Action-Objective” format.
+5.2 Main Results
+Table 2 presents the main results evaluated using the proposed dataset.
+Among the LLM-guided clustering baselines, the data-centric Keyphrase expansion method, which refines sentence embeddings via LLM-summarized keyphrases, achieved the most significant performance gain Viswanathan et al.
 (2024).
-Among the clustering algorithms, hierarchical clustering demonstrated the highest NMI and goodness scores, thus serving as the base model for the proposed methods.
-Sampling techniques (i.e., random and convex) are implemented to select representative sentences from each cluster as input for LLM.
-Among data-centric baselines, refining sentence embeddings through keyphrase expansion demonstrated the most improvement.
-However, deriving pairwise constraints for fine-tuning the embedding model, as proposed in ClusterLLM, proved ineffective due to the excessive number of clusters in the dataset.
-Interestingly, the evaluation of goodness does not always align with the NMI score, offering additional insights into cluster quality.
-For example, while keyphrase expansion improved the NMI score, the generated keyphrases could distort the meaning of the original sentence, resulting in less representative clusters and a slightly lower goodness score compared to the base model.
-The proposed LLM-in-the-loop intent clustering, abbreviated as the “LLM-ITL” approach, demonstrated satisfactory improvement in NMI and significant enhancement in the goodness score.
-In a context-free setting, the integration of LLM evaluation and naming outperformed both ClusterLLM and IDAS baselines.
-By incorporating effective keyphrase expansion, the NMI improved further, highlighting the potential of integrating data-centric methods into the pipeline.
-Notably, the proposed method does not require prior knowledge of the number of clusters and conducts parameter searches during the clustering process, whereas the baseline methods preset the number of clusters at $n=1507$.
-This adaptability underscores its significant application value for extracting intent clusters from any noisy text corpora Akama et al.
+However, deriving pairwise constraints for fine-tuning the embedding model, as in ClusterLLM, was ineffective due to the excessive number of clusters in the dataset.
+Interestingly, goodness evaluation does not always align with the NMI score.
+For example, while keyphrase expansion improved NMI, the generated keyphrases could distort the original sentence meaning, leading to less representative clusters and a slightly lower goodness score compared to the base model.
+This highlights the need to balance ground truth alignment with semantic coherence during evaluation.
+The proposed LLM-in-the-loop intent clustering demonstrated satisfactory improvements in NMI and significant enhancements in the goodness score.
+In a context-free setting, iterative clustering with convex sentence sampling outperformed both ClusterLLM and IDAS baselines.
+Effective keyphrase data augmentation further improved NMI, highlighting the potential of integrating data-centric methods into the pipeline.
+Furthermore, the context-aware approaches consistently improved performance by incorporating dialogue roles and cluster merging, resulting in a 4.48% increase in NMI without compromising the goodness score.
+This emphasizes the importance of task-centric design in leveraging intent labels effectively for the problem of intent clustering.
+Notably, the proposed method does not require prior knowledge of the number of clusters and performs parameter searches during clustering, unlike baseline methods that rely on the true number of clusters.
+This adaptability underscores its strong application potential for extracting intent clusters from noisy text corpora Akama et al.
 (2020).
-Furthermore, context-aware approaches were evaluated with the goal of enhancing the task performance of customer service dialogue intent clustering.
-A consistent improvement was observed when incorporating dialogue role into the clustering process, resulting in a 4.48% increase in NMI without compromising the goodness score.
-This highlights the importance of task-centric design in effectively addressing industrial problems.
+Method Bank77 CLINC(I) MTOP(I) Massive(I)
+SCCL 81.77 92.94 73.52 73.90
+Self-supervise 80.75 93.88 72.50 72.88
+ClusterLLM 85.07 94.00 73.83 77.64
+IDAS 82.84 92.35 72.31 75.74
+LLMEdgeRefine - 94.86 72.92 76.66
+ours 82.32 94.12 72.45 78.12
+Table 5: Performance (NMI) on English benchmarks.
+5.3 Evaluation on English Benchmarks
+To validate the effectiveness of the proposed LLM-in-the-loop technique, we evaluate the context-aware (best-performing) method on four widely used English benchmarks FitzGerald et al.
+(2023).
+The Llama-7b model is fine-tuned with 800 intent clusters (200 samples from each dataset) annotated by human experts to derive Action-Objective intent labels and bad clusters through perturbation.
+As shown in Table 5, our method delivers performance comparable to state-of-the-art LLM-guided methods and outperforms on Massive(I) dataset.
+Furthermore, the computational cost is lower than the compared methods as measured in Section 5.5.
 5.4 Application Performance
@@
 To demonstrate the practical advantages of the proposed methods, we further evaluate their performance by training BERT classifiers (bert-base-chinese model) based on the clustered results.
-Compared to the training data produced by the baseline method, which achieved only 65% classification accuracy, the proposed LLM-in-the-loop method improved the accuracy by 12%, providing promising evidence of the improved cluster quality.
+Compared to the training data produced by the base model, which achieved only 65% classification accuracy, the proposed LLM-in-the-loop method improved the accuracy by 12%, providing promising evidence of the improved cluster quality.
+5.5 Analysis of Computational Cost
+We systematically compare the cost of the proposed method with existing LLM-guided clustering on the Bank77 dataset, with a test set of $S=3,080$ sentences and true cluster number $N=77$.
+The proposed LLM-in-the-loop iterative method (Algorithm 1) uses candidate cluster numbers $N=[10,30,50,70]$, requiring $\sum_{n_{i}\in N}n_{i}=160$ LLM calls per iteration, and with $T=3$ iterations, totals $T\cdot\sum_{n_{i}\in N}n_{i}=480$ calls, each with 20 sentences input for coherence evaluation.
+ClusterLLM, employing a fixed triplet sampling approach with $Q=1,024$, incurs a constant $1,024$ LLM calls.
+Keyphrase expansion generates one keyphrase per sentence, resulting in a linear cost of $3,080$ LLM calls for the entire dataset.
+While our method requires more input tokens per LLM call, the cost remains efficient compared to its counterparts since the output token (i.e., “good/bad” label and intent label) is kept concise with fewer LLM calls.
 5.6 Ablation Studies
-We present two ablation studies to derive best practices for industrial deployment using different sampling and hyperparameter tuning strategies.
-6.4.1 Effective LLM Evaluation
-Optimizing the input to LLMs is crucial for enhancing performance and reducing costs Song et al.
+Effective Coherence Evaluation with Data Sampling and LLM Crowdsourcing
+Optimizing the input to LLM evaluator is crucial for enhancing performance and reducing costs Song et al.
 (2025); Jiang et al.
@@
 As shown in Figure 3, convex sampling outperforms random sampling in selecting representative sentences for coherence evaluation.
-Interestingly, as the sample time $t$ (for repeated evaluation) or the dimension of convex hull $d$ increases, performance consistently declines.
+Interestingly, as the sample time $t$ for repeated evaluation with majority voting or the dimension of the convex hull $d$ increases, performance consistently declines.
 This suggests that simpler, more efficient parameters lead to better accuracy.
 This pattern is verified through experiments with the ChatGLM model (see red line), which exhibits the same pattern.
+Figure 3: Performance comparison of sampling methods and hyperparameters for effective LLM evaluation.
 To better assess the consistency and alignment between LLM judgments and human perceptions, we utilize the concept of crowdsourcing for collaborative coherence evaluation Zhang et al.
@@
 The proposed LLM crowd consists of diverse entities Zhang et al.
-(2024) represented by various fine-tuned LLMs and uses a majority voting mechanism to aggregate their judgments Schoenegger et al.
+(2024) represented by four fine-tuned LLMs and uses a majority voting mechanism to aggregate their judgments Schoenegger et al.
 (2024).
 Based on the evaluation, we argue that a single fine-tuned LLM can effectively serve as a robust evaluator, as evidenced by the comparable performance to crowdsourced accuracy (cross markers in Figure 3) when using convex sampling.
-6.4.2 Parameter Search Space Pruning
-In practice, the search space for the number of clusters can be extensive, ranging from a few options to several hundred.
-To mitigate the computational costs associated with traditional model-based selection, we propose an LLM-native method for search space pruning, which predicts the optimal cluster number for subsequent iterations using logs from previous iterations (see Appendix 10).
-By pruning the search space and selecting the top 5 non-repetitive predicted solutions, we achieved an accuracy of 89.47% (see Table 5), indicating a high likelihood that the optimal cluster number is among these candidates.
-This approach substantially reduces redundant model fittings and enhances the efficiency of deploying the proposed methods.
+LLM qwen7b qwen14b baichuan2-7b chatglm3-6b
+Accuracy 62.50% 65.00% 58.50% 60.00%
+Table 6: Performance of fine-tuned LLMs in evaluating semantic coherence of intent clusters using numerical 1–4 scale scoring.
+Binary Judgment vs. Numerical Scoring
+To identify the optimal evaluation strategy for our clustering framework, we compared a numerical 1–4 scoring system (1: very poor, 4: very coherent) with the proposed binary good/bad judgment.
+We annotated 1,000 clusters with input from five human experts, finalized scores through majority voting, and fine-tuned four Chinese LLMs using this dataset.
+The models were then evaluated on 200 additional clusters, assigning scores five times per cluster and consolidating results via majority voting.
+The 1–4 scale demonstrated low accuracy and inconsistency, with 19% of samples observed to have significant disagreement (e.g., three ‘1’s and two ‘4’s) in Qwen-2.5-14B, making it unreliable for assessing cluster quality or aligning with human evaluation.
+In contrast, the simpler binary good/bad judgment provided consistent and interpretable quality assessments, proving to be the more effective evaluation strategy.
+Method NMI Goodness
+Geodesic distance (probabilistic) 0.8420 97.8%
+Cosine similarity (deterministic) 0.8102 94.7%
+Cosine similarity (probabilistic) 0.8242 95.2%
+Geodesic distance (deterministic) 0.8309 95.4%
+Table 7: Ablation study on cluster merging methods.
+Effectiveness of Cluster Merging Techniques
+Table 7 evaluates the impact of distance measures and merging decisions on the proposed cluster merging technique.
+Results highlight the superiority of geodesic distance in hyperspherical space, demonstrating its ability to effectively capture deeper semantic relationships compared to cosine similarity.
+Additionally, the probabilistic merging strategy consistently improves robustness by accounting for uncertainty, yielding better clustering quality regardless of the distance measure used.
+These findings emphasize the importance of both an appropriate distance metric and a probabilistic approach in achieving optimal merging outcomes.
 {CJK}
@@
 Table 8: Search space pruning by LLMs.
+Parameter Search Space Pruning
+In practice, the search space for the number of clusters can be extensive, ranging from a few options to several hundred.
+To mitigate the computational costs associated with traditional model-based selection, we propose an LLM-native method for search space pruning, which predicts the optimal cluster number for subsequent iterations using logs from previous iterations (see Appendix 13).
+By pruning the search space and selecting the top 5 non-repetitive predicted solutions, we achieved an accuracy of 89.47% (see Table 8), indicating a high likelihood that the optimal cluster number is among predicted candidates.
+This approach substantially reduces redundant model fittings and enhances the efficiency of deploying the proposed methods.
 6 Conclusion
-This paper tackles the dialogue intent clustering task by leveraging intermediate results through a human-aligned LLM-in-the-loop approach.
-Extensive experiments show that fine-tuned LLMs perform robustly in semantic coherence evaluations and cluster labeling.
-Both context-free and context-aware methods surpass existing LLM-guided clustering baselines, significantly improving cluster quality and downstream application performance.
-The ablation studies further provide practical insights for efficient industrial deployment.
-Future research is encouraged to enhance LLM reliability by integrating confidence estimation into the evaluation process, as well as employing advanced base algorithms and more sophisticated judgment formats to reveal additional insights.
+This paper tackles dialogue intent clustering using a human-aligned LLM-in-the-loop approach.
+Experiments on a complex, realistic Chinese customer service dataset reveal that fine-tuned LLM utilities excel in semantic coherence evaluation and cluster labeling, with the proposed methods consistently outperforming existing LLM-guided clustering baselines in cluster quality and computational efficiency.
+Future work should focus on refining evaluation strategies to enhance semantic guidance from LLMs and exploring task-centric approaches for deeper LLM-in-the-loop integration.
+Limitations
+This work has two main limitations.
+First, while the coherence evaluation is practical and intuitive, relying solely on cluster coherence as the quality indicator overlooks other critical attributes, such as meaningfulness and expressiveness, which are equally important in assessing a cluster’s usefulness.
+For instance, a cluster labeled “express-feeling” may be overly general and should be refined into more specific categories like “express-appreciation” and “state-dissatisfaction” to enhance interpretability and practical applicability.
+Additionally, the binary evaluation and numerical scoring considered in this paper are deterministic indicators, overlooking probabilistic judgments and confidence intervals, which could provide deeper insights into clustering assessment and offer greater flexibility for integration into the LLM-in-the-loop clustering process.
+Second, the dataset proposed in this study is limited to the Chinese language and focuses on specific domains, namely banking, insurance, and telecommunication.
+While it represents the largest realistic customer service call dataset for the intent clustering problem, its coverage is restricted in terms of multi-domain and multilingual aspects, potentially affecting the generalizability of our findings to other languages and domains.
+Furthermore, the scope of this paper is confined to intent clustering, which serves as the initial step in the broader process of intent discovery.
+Subsequent steps, such as analyzing intent trajectories and recognizing intents at the document level rather than the sentence level, are not comprehensively discussed.
+These areas warrant further investigation to fully reveal the practical values of LLM-in-the-loop design in intent mining tasks.
 Appendix A Data Collection and Annotation
@@
 The original dialogue dataset is sourced from audio transcriptions of customer service calls across several domains.
-Initially, the dataset contains over 120,000 calls.
+Initially, the dataset contains over 11,879 customer service calls.
 Strict filtering is applied to maintain data integrity and exclude confidential or sensitive content, resulting in 8,184 dialogues and 69,839 sentences.
@@
 Data Annotation.
-To create the intent clustering dataset, we recruited 15 human experts, each with over five years of experience in customer service call centers, to participate in the data annotation process:
+To construct the intent clustering dataset, we internally recruited 15 human experts, each with over five years of experience in customer service call centers, to participate in the data annotation process:
   1. 1.
@@
 The final dataset consists of 1,507 high-quality intent clusters.
-Appendix B Prompt Template and More Results for LLM Utilities
-The prompt for coherence evaluation and cluster naming are translated into English for demonstration.
-For each input, a few-shot demonstration with five input-output pairs is provided to ensure the consistency of the output format and enhance understanding of the task.
+Appendix B Prompt Template
+The prompts for coherence evaluation and cluster naming are translated into English for demonstration purposes.
+Each input is accompanied by a few-shot demonstration with five input-output pairs to ensure consistency in the output format and enhance understanding of the task.
 Coherence Evaluation -  Your are a helpful assistant for sentence clustering.
@@
 Example: input:[sentences] output:[label] input:{[sentences]} output:
-Considering the resource constraints of industrial deployment, we evaluated the coherence performance using the OpenAI APIs: GPT-3.5, GPT-4, and GPT-4o, achieving accuracies of 89.58%, 93.54%, and 94.17%, respectively.
-These findings indicate that a smaller, fine-tuned LLM, such as qwen7b, can outperform these advanced models and potentially reduce the costs associated with API requests [^3] The fine-tuned LLMs for coherence evaluation and cluster naming will be released upon acceptance..
-To highlight the importance of fine-tuning, we also tested a vanilla Qwen2.5 model without fine-tuning, which resulted in a significantly lower accuracy of 75.63%.
-Model Sampling Method 1 2 3 4 5
-qwen14b convex - 96.25% 95.42% 95.42% 95.21%
-random (n=10) 95.83% 94.38% 93.33% 92.92% 92.92%
-random (n=20) 94.58% 95.00% 94.38% 94.17% 94.17%
-chatglm3-6b convex - 92.08% 90.83% 91.04% 90.42%
-random (n=10) 92.29% 88.75% 85.83% 84.58% 85.21%
-random (n=20) 90.42% 89.58% 89.38% 89.58% 88.98%
-Figure 2: Performance comparison of sampling methods and hyperparameters for effective LLM evaluation.
+Appendix C More Results: Evaluation with Proprietary LLM
+Given the resource-intensive nature of LLM fine-tuning, we conducted further experiments in coherence evaluation using proprietary LLMs accessed through OpenAI APIs: GPT-3.5, GPT-4, and GPT-4o.
+These advanced and widely used models offer strong performance in various tasks but also come with higher operational costs due to token consumption.
+Using the same evaluation set of 480 clusters, these models achieved accuracies of 89.58%, 93.54%, and 94.17%, respectively.
+The results highlight that a smaller, fine-tuned LLM, such as Qwen-2.5-7b (with an accuracy of 96.25%), can outperform these advanced models while significantly reducing costs associated with API usage [^4] The fine-tuned Chinese LLMs for coherence evaluation and cluster naming will be released upon acceptance..
+For comparison, a vanilla Qwen-2.5-7b model (without fine-tuning) was also evaluated, yielding a substantially lower accuracy of 75.63%, emphasizing the essence of fine-tuning.
 {CJK}

### v4

@@
 Abstract
-Discovering customer intentions in dialogue conversations is crucial for automated service agents.
-However, existing intent clustering methods often fail to align with human perceptions due to a heavy reliance on embedding distance metrics and a tendency to overlook underlying semantic structures.
-This paper proposes an LLM-in-the-loop (LLM-ITL) intent clustering framework, integrating the semantic understanding capabilities of LLMs into conventional clustering algorithms.
-Specifically, this paper (1) investigates the effectiveness of fine-tuned LLMs in semantic coherence evaluation and intent cluster naming, achieving over 95% accuracy aligned with human judgments; (2) designs an LLM-ITL framework that facilitates the iterative discovery of coherent intent clusters and the optimal number of clusters; and (3) proposes context-aware techniques tailored for customer service dialogue.
-As existing English benchmarks offer limited semantic diversity and intent groups, we introduce a comprehensive Chinese dialogue intent dataset, comprising over 100k real customer service calls and 1,507 human-annotated intent clusters.
-The proposed approaches significantly outperform LLM-guided baselines, achieving notable enhancements in clustering quality and lower computational cost.
-Combined with several best practices, our findings highlight the potential of LLM-in-the-loop techniques for scalable and human-aligned intent clustering.
+Discovering customer intentions is crucial for automated service agents, yet existing intent clustering methods often fall short due to their reliance on embedding distance metrics and neglect of underlying semantic structures.
+To address these limitations, we propose an LLM-in-the-loop (LLM-ITL) intent clustering framework, integrating the language understanding capabilities of LLMs into conventional clustering algorithms.
+Specifically, this paper (1) examines the effectiveness of fine-tuned LLMs in semantic coherence evaluation and intent cluster naming, achieving over 95% accuracy aligned with human judgments; (2) designs an LLM-ITL framework that facilitates the iterative discovery of coherent intent clusters and the optimal number of clusters; and (3) introduces context-aware techniques tailored for customer service dialogue.
+Since existing English benchmarks lack sufficient semantic diversity and intent coverage, we further present a comprehensive Chinese dialogue intent dataset comprising over 100k real customer service calls with 1,507 human-annotated clusters.
+The proposed approaches significantly outperform LLM-guided baselines, achieving notable improvements in clustering quality, cost efficiency, and downstream applications.
+Combined with several best practices, our findings highlight the prominence of LLM-in-the-loop techniques for scalable dialogue data mining.
 Dial-In LLM: Human-Aligned LLM-in-the-loop Intent Clustering for Customer Service Dialogues
-Mengze Hong1, Wailing Ng1, Chen Jason Zhang1 Yuanfeng Song2, Di Jiang2thanks: Corresponding author: dijiang@webank.com 1Hong Kong Polytechnic University 2AI Group, WeBank Co., Ltd
+Mengze Hong1thanks: Work was partially done during internship at WeBank., Wailing Ng1, Chen Jason Zhang1, Yuanfeng Song2, Di Jiang1thanks: Corresponding Author 1Hong Kong Polytechnic University, 2AI Group, WeBank Co., Ltd
 1 Introduction
 Figure 1: Comparison of embedding-distance metric and human perception: cosine similarity failed to identify the same intentions under diverse expression (top), and to distinguish distinct intentions under similar expressions (bottom).
-Intent discovery is a crucial task in numerous NLP applications, including the design of AI-driven dialogue systems Hengst et al.
-(2024) and the learning of utterance patterns Ghosal et al.
+Intent discovery is a crucial task in NLP applications, such as dialogue system design Hengst et al.
+(2024), information retrieval Jiang et al.
+(2016a), and utterance pattern analysis Ghosal et al.
 (2020).
@@
 (2023), existing research primarily focuses on developing meaningful sentence representations and relies on embedding distance metrics for optimization Yin et al.
+(2021); Gao et al.
 (2025).
-This approach often overlooks the distinctive characteristics of textual information, such as linguistic patterns and semantic diversity (see Figure 1), and restricts human-aligned evaluation and validation of clustering quality Vinh et al.
+This approach often overlooks the distinctive characteristics of textual information, such as linguistic patterns and semantic diversity (see Figure 1), and thus restricts human-aligned evaluation and validation of clustering performance Vinh et al.
 (2009).
-This issue is particularly pronounced in languages with rich semantics, such as Chinese, where two seemingly similar sentences can convey entirely different meanings, and vice versa Chen (1993).
+This issue is particularly pronounced in languages with rich semantics, such as Chinese, where two seemingly similar sentences can convey entirely different meanings Chen (1993); Jiang et al.
+(2016b).
 With growing research interest in integrating large language models (LLMs) into the problem-solving pipeline Hong et al.
-(2025), LLM-guided clustering techniques have emerged, demonstrating superior performance over traditional machine learning algorithms.
-While these methods effectively incorporate language understanding into the clustering process, they primarily focus on data preprocessing, querying LLMs for embedding refinement or data augmentation Zhang et al.
-(2023); Viswanathan et al.
+(2025b), LLM-guided clustering techniques have emerged, demonstrating superior performance over traditional machine learning algorithms Zhang et al.
 (2023).
-This approach represents a surface-level integration, potentially missing the benefits of LLMs in providing semantic-driven guidance to the clustering process and the quality of the final results.
-In this paper, we introduce an LLM-in-the-loop (LLM-ITL) intent clustering method, designed to facilitate the iterative discovery of coherent intent clusters.
-Our approach adeptly utilizes intermediate clustering results by incorporating human-aligned LLM utilities, offering computationally efficient and human-interpretable intent clustering.
+While these methods effectively incorporate language understanding into the clustering process, they primarily focus on data preprocessing, querying LLMs for embedding refinement or data augmentation Viswanathan et al.
+(2024).
+This approach represents a surface-level integration, potentially missing the benefits of LLMs to contribute semantic-driven guidance within the clustering process.
+In this paper, we introduce an LLM-in-the-loop (LLM-ITL) intent clustering framework, designed to facilitate the iterative discovery of coherent intent clusters from semantically diverse, large-scale dialogue datasets.
+Our approach effectively leverages intermediate clustering results by incorporating human-aligned LLM utilities, enabling computationally efficient and human-interpretable intent clustering.
 The key contributions of this work are:
   1. 1.
-    We introduce the largest Chinese dialogue intent clustering dataset, notable for its high semantic diversity and a significant inclusion of noisy, out-of-domain intents.
-This dataset is derived from over 100,000 real-world customer service call sessions, comprising 1,507 intent clusters annotated by human experts.
-It fills a crucial gap in evaluation benchmarks for complex intent clustering and classification.
+    We present the largest Chinese dialogue intent clustering dataset, derived from over 100,000 real-world customer service calls across the banking, telecommunication, and insurance domains.
+The data is annotated into 1,507 intent clusters with high semantic diversity and a substantial inclusion of noisy, out-of-domain queries, reflecting realistic and complex customer interactions.
   2. 2.
-    We demonstrate the effectiveness of fine-tuned LLMs in assessing the semantic coherence of intent clusters and accurately providing intent labels across various sampling strategies.
-This highlights two effective and efficient utilities for designing LLM-in-the-loop solutions tailored for intent clustering.
+    We demonstrate the effectiveness of fine-tuned small LLMs in assessing the semantic coherence of intent clusters and providing accurate intent labels across various sampling strategies, offering cost-efficient utilities for designing LLM-in-the-loop solutions.
   3. 3.
-    We propose an LLM-in-the-loop clustering algorithm to derive an optimal set of clusters in scenarios where the number of clusters is unknown.
-This approach effectively leverages intermediate results generated during the clustering process, outperforming state-of-the-art baselines and providing practical benefits for downstream applications.
-Additionally, the discussion on data sampling and LLM-based crowdsourcing further improves performance.
+    We propose an LLM-in-the-loop intent clustering framework that effectively combines the strengths of LLMs and conventional clustering algorithms.
+This approach outperforms state-of-the-art baselines and excels in downstream applications with 18.46% performance gain.
+Furthermore, discussions on data sampling and LLM-based crowdsourcing validate best practices for real-world deployment.
+Figure 2: Overview of the proposed LLM-in-the-loop framework for dialogue intent clustering.
 2 Related Work
-This study is grounded in two key areas of research: LLM-guided text clustering and intent clustering.
 LLM-guided Text Clustering.
@@
 Recently, Hong et al.
-(2025) conceptualized LLM-in-the-loop machine learning, classifying LLM integration into data-, model-, and task-centric strategies.
-This method mirrors a “human-in-the-loop” approach Chen et al.
-(2024), facilitating effective LLM integration in machine learning applications.
+(2025b) introduced the concept of LLM-in-the-loop machine learning, categorizing integration strategies into data-, model-, and task-level approaches.
+This paradigm parallels the human-in-the-loop framework Chen et al.
+(2024), enabling LLMs to replicate human expertise and thereby enhance conventional problem-solving workflows in a cost-efficient and flexible manner.
+Within this taxonomy, existing work on LLM-guided clustering has primarily focused on data-centric aspects, with relatively limited exploration of modeling and task-solving aspects.
+This gap constrains a broader understanding of the potential of LLMs in tackling long-standing challenges in text clustering, such as improving cluster interpretability and moderating the clustering process Tan et al.
+(1999); Jiang et al.
+(2021).
 Intent Clustering.
-Intent clustering is a fundamental step in intent induction Chandrakala et al.
-(2024) and the development of intent classifiers Gung et al.
+Intent clustering extends conventional text clustering by incorporating contextual cues and domain knowledge to uncover meaningful user intentions Allahyari et al.
+(2017); Hong et al.
 (2024).
-While semi-supervised Kumar et al.
+It serves as a fundamental step in intent induction Chandrakala et al.
+(2024) and in preparing data for training intent classifiers Gung et al.
+(2023).
+While conceptually related to topic modeling Jiang et al.
+(2015, 2023), which seeks to uncover latent semantic themes, intent clustering tackles the more challenging task of differentiating texts that may appear lexically similar but convey contextually distinct meanings, thereby uncovering the underlying communicative goals behind sentences Carberry and Flowers (1988); Qu et al.
+(2018).
+This finer granularity makes intent clustering more demanding as a data mining task and a critical component in applications such as search engines Jiang et al.
+(2013, 2016a) and dialogue systems Qin et al.
+(2023); Hong et al.
+(2025c).
+While semi-supervised approaches Kumar et al.
 (2022) and deep learning methods Lin et al.
 (2020); Zhang et al.
-(2021) have significantly advanced intent clustering, recent progress in LLM-based systems has highlighted their potential practical advantages Liang et al.
+(2021) have substantially advanced intent clustering, recent attempts in LLM-based systems highlight their practical advantages Liang et al.
 (2024).
-The IDAS method effectively employs in-context learning to generate descriptive utterance labels and enhance sentence embeddings De Raedt et al.
+For instance, the IDAS method leverages in-context learning to generate descriptive utterance labels and improve sentence embeddings De Raedt et al.
 (2023).
-However, challenges such as computational cost and robustness remain unresolved Song et al.
-(2023).
+However, challenges such as high computational cost and limited model robustness remain unresolved Song et al.
+(2023), and existing methods are evaluated based on small-scale datasets with limited complexity Casanueva et al.
+(2020), leading to an inadequate understanding of their effectiveness in real-world applications.
 Highlights.
 Previous studies have primarily focused on enhancing input text representations and the direct use of LLMs for cluster modification, imposing high computational cost and uncertainty.
-In contrast, this paper implements human-aligned LLM utilities to create a LLM-in-the-loop intent clustering process with emphasis on the intermediate results.
-Additionally, this work tackles a more realistic scenario by releasing a complex dataset transcribed from real customer service calls, offering practical insights and contributing to more meaningful evaluations in future intent clustering.
-Figure 2: Overview of proposed LLM-in-the-loop Dialogue Intent Clustering.
+In contrast, this paper implements human-aligned LLM utilities to create an LLM-in-the-loop framework with emphasis on the intermediate clustering results.
+Additionally, this work releases a complex dataset from real customer service call transcriptions, enabling more practical insights and meaningful evaluations for future intent clustering.
 3 Proposed Methods
@@
 Coherence Evaluation.
-Inspired by human behaviors in perceiving texts and making comparisons based on semantic meaning rather than word-level similarity Peter W.
-Foltz and Landauer (1998), this paper proposes the semantic consistency within a cluster, known as semantic coherence, as a superior metric and optimization target.
-The coherence evaluation is formulated as a binary classification problem, leading toward intuitive evaluation and easing the training of the LLM evaluator (see Section 5.6 for different formulations).
-Good clusters consist of sentences focused on a specific topic, while Bad clusters contain inconsistent or ambiguous intentions (see Table 9 for examples).
-This task is particularly challenging for traditional machine learning classifiers due to a lack of semantic understanding Mimno et al.
-(2011), necessitating fine-tuned LLMs Gu et al.
-(2025) for robust evaluation (see Section 5.1 for fine-tuning details).
+Inspired by human behaviors in perceiving texts and making comparisons based on semantic meaning rather than surface-level work similarity Peter W.
+Foltz and Landauer (1998), this paper proposes semantic coherence as a more effective metric and optimization objective, measuring the semantic consistency within a cluster.
+The coherence evaluation is formulated as a binary classification problem, enabling intuitive interpretation and simplifying the training of the LLM evaluator (see Section 5.6 for alternative formulations).
+Good clusters consist of sentences focused on a specific topic, whereas Bad clusters contain inconsistent or ambiguous intentions (see Table 11 for examples).
+This task is particularly challenging for traditional machine learning classifiers due to their lack of semantic understanding Mimno et al.
+(2011), thus necessitating fine-tuned LLMs for robust evaluation Gu et al.
+(2025).
 Cluster Naming.
 Giving each cluster a concise and meaningful name is essential for many downstream applications Pattnaik et al.
+(2024); Luo et al.
 (2024).
-In this paper, we introduce a novel naming convention, "Action-Objective," which is particularly effective for capturing dialogue intents that are typically topic-oriented (e.g., insurance, loan) and involve distinct actions (e.g., inquire, confirm).
-Examples of human annotations are shown in Table 10, and a comparison of different naming conventions is presented in Table 11, demonstrating the superiority of the proposed convention in capturing customer intention with clear goal-oriented expression.
+In this paper, we introduce a novel naming convention, ”Action-Objective,” which is particularly effective for capturing dialogue intents that are typically topic-oriented (e.g., insurance, loan) and involve distinct actions (e.g., inquire, confirm).
+Examples of human annotations are shown in Table 12, and a comparison of different naming conventions is presented in Table 13 to demonstrate the effectiveness of the proposed approach.
 3.2 LLM-in-the-loop Iterative Intent Clustering with Coherence Evaluation
+At iteration $t$, the current set of unassigned sentences is denoted as $\mathcal{S}^{(t)}$.
+A special case in the first iteration, where $\mathcal{S}^{(0)}=\mathcal{S}$ represents the entire set of unique sentences derived from the dialogue corpus.
+For each candidate cluster number $n_{i}\in N$, we compute:
+$\mathcal{C}_{n_{i}}^{(t)}=F(\mathcal{E}^{(t)},n_{i}),$
+where $F$ is a clustering function (e.g., K-means clustering), and $\mathcal{E}^{(t)}=\{\mathbf{e}_{s}\mid s\in\mathcal{S}^{(t)}\}$ represents the sentence embeddings.
+This results in $|N|$ distinct cluster assignments as the initial outcome.
+Then, the semantic coherence of each cluster is evaluated using a fine-tuned LLM $\mathcal{M}_{\text{eval}}$:
+$\mathbf{g}_{n_{i}}^{(t)}=\left[\mathcal{M}_{\text{eval}}(C_{1}^{(t)}),\ldots,\mathcal{M}_{\text{eval}}(C_{n_{i}}^{(t)})\right],$
+where $\mathcal{M}_{\text{eval}}(C_{k})=1$ if the cluster is coherent (i.e., “good” cluster), else 0.
+While the number of clusters is often known in benchmark evaluations, determining the optimal number in a noisy text corpus is both challenging and essential.
+Here, we propose a local search heuristic that maximizes the “good/bad” ratio at each iteration.
+The optimal $n_{t}^{*}$ is given by:
+$n_{t}^{*}=\arg\max_{n_{i}\in N}\frac{\sum_{j=1}^{n_{i}}\mathbb{I}\left(\mathbf{g}_{n_{i}}^{(t)}[j]=1\right)}{\sum_{j=1}^{n_{i}}\mathbb{I}\left(\mathbf{g}_{n_{i}}^{(t)}[j]=0\right)+1},$
+representing the best cluster number at iteration $t$.
+This approach enables the automatic discovery of suitable cluster numbers in a step-by-step manner.
+The search space $\bm{N}$ should be selected carefully to balance accuracy and efficiency, and a search space pruning strategy is proposed in Section 5.6 to enhance the searching process.
+Finally, the “good” clusters in the optimal $\mathcal{C}_{n^{*}}^{(t)}$ are retained, and the remaining sentences will be refined in the next (i.e., $t+1$) iteration, enabling the iterative discovery of high-quality clusters.
+The proposed method is summarized in Algorithm 1, and the LLM integration is highlighted.
 Algorithm 1 LLM-in-the-loop Intent Clustering
@@
 2:Set of clusters $\mathcal{C}$
-3:$\mathcal{E}\leftarrow\{f_{\text{emb}}(s)\mid s\in\mathcal{S}\}$
-4:$\mathcal{S}^{(0)}\leftarrow\mathcal{S}$, $\mathcal{C}\leftarrow\emptyset$, $t\leftarrow 0$
+[…]
+[…]
 5:while $\frac{|\mathcal{S}^{(t)}|}{|\mathcal{S}|}>\epsilon$ and $t<T_{\text{max}}$ do
@@
 17:return $(\mathcal{C})$
-At iteration $t$, the current set of unassigned sentences is denoted as $\mathcal{S}^{(t)}$.
-A special case in the first iteration, where $\mathcal{S}^{(0)}=\mathcal{S}$ represents the entire set of unique sentences derived from the dialogue corpus.
-For each candidate cluster number $n_{i}\in N$, we compute:
-$\mathcal{C}_{n_{i}}^{(t)}=F(\mathcal{E}^{(t)},n_{i}),$
-where $F$ is a clustering function (e.g., K-means clustering), and $\mathcal{E}^{(t)}=\{\mathbf{e}_{s}\mid s\in\mathcal{S}^{(t)}\}$ representing the sentence embeddings.
-This results in $N$ distinct cluster assignments.
-Then, the semantic coherence of each cluster is evaluated using a fine-tuned LLM $\mathcal{M}_{\text{eval}}$:
-$\mathbf{g}_{n_{i}}^{(t)}=\left[\mathcal{M}_{\text{eval}}(C_{1}^{(t)}),\ldots,\mathcal{M}_{\text{eval}}(C_{n_{i}}^{(t)})\right],$
-where $\mathcal{M}_{\text{eval}}(C_{k})=1$ if the cluster is coherent (i.e., “good” cluster), else 0.
-While the number of clusters is often known with benchmark evaluation, it is challenging but necessary to implement mechanisms for determining the optimal number of clusters in a noisy text corpus.
-Here, we propose a local search heuristic that maximizes the “good/bad” ratio at each iteration.
-The optimal $n_{t}^{*}$ is selected as:
-$n_{t}^{*}=\arg\max_{n_{i}\in N}\frac{\sum_{j=1}^{n_{i}}\mathbb{I}\left(\mathbf{g}_{n_{i}}^{(t)}[j]=1\right)}{\sum_{j=1}^{n_{i}}\mathbb{I}\left(\mathbf{g}_{n_{i}}^{(t)}[j]=0\right)+1},$
-where $n_{t}$ denotes the best cluster number at iteration $t$.
-This enables the automatic discovery of suitable cluster numbers in a step-by-step manner.
-The search space $\bm{N}$ should be chosen carefully to balance accuracy and efficiency; thus, we propose a search space pruning strategy, as illustrated in Section 5.6 for more details.
-Finally, the “good” clusters in the optimal $\mathcal{C}_{n^{*}}^{(t)}$ are retained, and the remaining sentences will be refined in the next (i.e., $t+1$) iteration, enabling the iterative discovery of high-quality clusters.
-The proposed method is summarized in Algorithm 1, and the LLM integration is highlighted.
 3.3 Post-Correction with LLM-Generated Intent Labels
-Preliminary results in Table 12 suggest that multiple clusters may capture similar intents due to diminishing size of unassigned sentences $\mathcal{S}^{(t)}$ in later iterations, leading to smaller and less representative clusters.
+Preliminary results in Table 14 suggest that, in later iterations, the diminishing size of the unassigned sentence set $\mathcal{S}^{(t)}$ may lead to the formation of multiple clusters capturing similar intents, resulting in smaller and less representative clusters.
 The embedding distances between sentences within clusters limit the natural consolidation of clusters with similar intents but different expressions Khan et al.
 (2020), necessitating a post-correction step to merge semantically aligned clusters.
-In contrast to previous methods that rely on direct LLM queries for validating individual cluster assignments Viswanathan et al.
+Previous methods typically address this by issuing direct LLM queries to validate individual cluster assignments Viswanathan et al.
 (2024); Feng et al.
-(2024), which are computationally expensive and inconsistent, we propose a context-aware approach, leveraging LLMs’ naming utility to robustly merge clusters based on intent labels.
+(2024), an approach that is both computationally expensive and prone to inconsistency.
+In contrast, we propose a context-aware approach, leveraging LLMs’ naming utility to robustly merge clusters based on their generated intent labels.
 At the end of the iterative clustering process, each cluster $C_{k}$ receives an intent label from the fine-tuned LLM naming utility:
@@
 with $\pi_{m}=1/K$ as uniform mixture weights.
-An edge is retained only if this probability exceeds a predefined threshold $\tau$ ($\tau=0.7$), minimizing inappropriate merges by accounting for uncertainty.
+An edge is retained only if this probability exceeds a predefined threshold $\tau$ (e.g., $\tau=0.7$), minimizing inappropriate merges by accounting for uncertainty.
 Finally, clusters linked in the affinity graph are consolidated into connected components, forming a refined cluster assignment $\mathcal{C}^{\prime}=\{C^{\prime}_{1},C^{\prime}_{2},\ldots,C^{\prime}_{K^{\prime}}\}$ that eliminates redundancy, enhances interpretability, and maintains robust semantic alignment.
@@
 3.4 Context-Aware Role Separation with LLM-Generated Intent Labels
-From a practical perspective, dialogue datasets are often accompanied by domain-specific features or use-case scenarios (e.g., customer service calls, group discussions).
-This motivates the incorporation of contextual information into the clustering process.
-The customer service dialogues typically involve only the customer and the service agent Lin et al.
+From a practical perspective, dialogues are often accompanied by domain-specific features or use-case scenarios (e.g., customer service calls, group discussions).
+This motivates the incorporation of contextual information into the clustering process Ding et al.
+(2025).
+In particular, the customer service dialogues typically involve only the customer and the service agent Lin et al.
 (2022).
 The classification task that assigns a sentence to its associated role is relatively simple with labeled training data.
-To provide an unsupervised solution, the sentence roles are naturally determined based on the LLM- intent label.
+To provide an unsupervised solution, we propose that the sentence roles can be naturally determined based on the LLM-generated intent label.
 By heuristics, sentences within clusters labeled as “inquire-” or similar actions are mostly sent from the customer, and vice versa.
 A two-step approach is proposed: initially, intermediate clustering results are obtained using previous approach, denoted as $C_{\text{inter}}=\{C_{1},C_{2},\ldots,C_{K}\}$, and these results are divided based on intent labels $l_{k}$, forming two groups with distinct roles, $R_{\text{customer}}$ and $R_{\text{agent}}$, where:
-$\begin{array}[]{c}R_{\text{customer}}=\{s\in S\mid l_{k}(s)\in\{\text{"inquire-"},\dots\}\},\\ R_{\text{agent}}=\{s\in S\mid l_{k}(s)\notin\{\text{"answer-"},\dots\}\}.\end{array}$
-In the second step, the sentences corresponding to each role are clustered again, denoted as $C^{\prime}_{\text{customer}}=\text{Cluster}(R_{\text{customer}})$ and $C^{\prime}_{\text{agent}}=\text{Cluster}(R_{\text{agent}})$, to achieve more refined outcomes, allowing for a clear separation of customer intent from service agent intent.
-Finally, the two groups of clusters are merged together, serving as the final or intermediate result to be further processed.
+$\begin{array}[]{c}R_{\text{customer}}=\{s\in S\mid l_{k}(s)\in\{\text{"inquire-"},\dots\}\},\\ R_{\text{agent}}=\{s\in S\mid l_{k}(s)\in\{\text{"answer-"},\dots\}\}.\end{array}$
+In the second step, the sentences corresponding to each role are clustered again, denoted as $C^{\prime}_{\text{customer}}=\text{Cluster}(R_{\text{customer}})$ and $C^{\prime}_{\text{agent}}=\text{Cluster}(R_{\text{agent}})$.
+Finally, the resulting customer and agent clusters are merged to produce a refined clustering that maintains a clear separation of intents by role, serving either as the final output or as an improved intermediate stage for further processing.
 4 Experiment
 4.1 Dataset
-The proposed dataset contains 1,507 high-quality intent clusters manually annotated from over 100,000 realistic customer service calls, comprising 55,085 distinct sentences with an average length of 17 Chinese characters per sentence (see Appendix A for annotation details) [^1] The full dataset and fine-tuned models will be released upon acceptance.
-Sample dataset can be found at GitHub..
-Among these intents, 885 are identified as domain-specific queries (e.g., inquire-insurance), while 622 are categorized as out-of-domain (e.g., provide-location), representing general queries that commonly arise in a customer service interaction.
-This dataset is the first Chinese benchmark for intent clustering and the largest of its kind compared to the existing English benchmark datasets shown in Table 1.
+The proposed dataset contains 1,507 high-quality intent clusters manually annotated from over 100,000 realistic customer service calls, comprising 55,085 distinct sentences with an average length of 17 Chinese characters per sentence (see Appendix A for annotation details) [^1] Data is available at GitHub repository..
+Among these intents, 885 are identified as domain-specific (e.g., inquire-insurance), primarily concentrated within the banking, telecommunications, and insurance industries, with a focus on the Chinese context Hong et al.
+(2025a).
+The remaining 622 clusters are categorized as out-of-domain (e.g., provide-location), representing general queries that commonly arise in customer service interactions.
+Compared with existing intent clustering benchmarks such as BANK77 Casanueva et al.
+(2020) and CLINC150 Larson et al.
+(2019), the proposed dataset is the first Chinese benchmark for customer service intent clustering and the largest of its kind in both the number of sentences and the number of clusters (see Table 1).
 It presents several new challenges, including the lexical sparsity in short-text sentences, the dependence on contextual knowledge, and the difficulties in balancing accuracy with computational efficiency due to the excessively large intent size.
 The semantic diversity, calculated based on the average cosine distance between individual sentences and the cluster centroid Casanueva et al.
-(2022), depicts the complexity of this dataset and motivates the development of semantic-guided techniques.
+(2022), depicts the complexity of this dataset and motivates the development of semantic-guided approaches.
 Number of Number of Semantic
@@
 BANK77 3080 77 0.209
-OOS 23,700 150 0.254
 NLU++ 3,080 62 0.367
@@
 ours 55,085 1507 0.538
-Table 1: Comparison of proposed dataset with existing intent clustering datasets.
-Category Method NMI #Good
-Baselines K-Means 0.7899 94.8%
-GMMs 0.7903 91.1%
-Hierarchical* 0.8001 94.9%
-LLM-Guided Clustering Baselines ClusterLLM Zhang et al. (2023) 0.7284 91.2%
-IDAS De Raedt et al. (2023) 0.8109 93.6%
-Keyphrase Viswanathan et al. (2024) 0.8371 94.5%
-LLMEdgeRefine Feng et al. (2024) 0.7411 87.2%
-Proposed Method (context-free) LLM-ITL (random) 0.8202 97.7%
-LLM-ITL (convex) 0.8208 97.8%
-LLM-ITL + keyphrase 0.8378 96.4%
-Proposed Method (context-aware) LLM-ITL + merge 0.8420 97.8%
-LLM-ITL + role 0.8679 97.2%
-LLM-ITL + role + merge 0.8826 97.6%
-Table 2: Performance of baselines and proposed methods in dialogue intent clustering. The best results are bolded.
-{CJK}
-UTF8gkai
-LLM qwen7b qwen14b baichuan2-7b chatglm3-6b
-Accuracy 96.25% 97.50% 89.17% 95.83%
-Table 3: Performance of fine-tuned LLMs in evaluating semantic coherence of intent clusters.
-{CJK}
-UTF8gkai
-LLM qwen7b qwen14b baichuan2-7b chatglm3-6b
-Accuracy 92.8% 94.3% 94.3% 94.4%
-Table 4: Performance of fine-tuned LLMs in naming intent clusters.
+Table 1: Comparison of the proposed customer service intent clustering dataset with existing benchmarks.
 4.2 Metrics
-Normalized Mutual Information (NMI) measures the degree of similarity between ground-truth and predicted clusters, scaled between 0 (no mutual information) and 1 (perfect correlation).
-However, researchers have claimed biased behavior in NMI Jerdee et al.
-(2024), necessitating further evaluation that reflects the true quality of the intent cluster from human perception.
+Normalized Mutual Information (NMI) measures the degree of similarity between ground-truth and predicted clusters, ranging from 0 (no mutual information) to 1 (perfect correlation).
+However, recent work has demonstrated that NMI exhibits biased behavior, particularly in favor of larger numbers of clusters Jerdee et al.
+(2024), highlighting the need for complementary evaluation metrics that better reflect human-perceived clustering quality.
 Goodness score measures the proportion of good clusters evaluated by the fine-tuned LLM evaluator.
@@
 For coherence evaluation, a human-annotated training dataset of 1,772 intent clusters labeled as “good” or “bad” is used.
-For cluster naming, a training dataset of 2,500 clusters, each containing 20 sentences, is annotated with intent labels.
+For cluster naming, a training dataset of 2,500 clusters, each containing 20 sentences, is annotated with the “Action-Objective” labels.
 Semantic-rich embeddings are generated using the BAAI General Embeddings (BGE) model [^3] https://huggingface.co/BAAI/bge-large-zh-v1.5 Xiao et al.
@@
 The best-performing clustering algorithm, hierarchical clustering, and LLM utilities, qwen14b and chatglm3-6b, are selected for LLM-in-the-loop clustering.
-To reduce computational load, both random and convex sampling methods are implemented to select 20 representative sentences per cluster as input for the LLM.
+Additionally, we employ random and convex sampling to select 20 representative sentences per cluster as LLM inputs, reducing computational overhead while preserving intent coverage.
 5 Results and Discussions
+LLM
+qwen7b
+qwen14b
+baichuan2-7b
+chatglm3-6b
+Accuracy 96.25% 97.50% 89.17% 95.83%
+Table 2: Performance of fine-tuned LLMs in evaluating semantic coherence of intent clusters.
+LLM
+qwen7b
+qwen14b
+baichuan2-7b
+chatglm3-6b
+Accuracy 92.8% 94.3% 94.3% 94.4%
+Table 3: Performance of fine-tuned LLMs in naming intent clusters.
 5.1 LLM Utilities
@@
 The results in Table 3 demonstrate that the fine-tuned LLMs show promising performance in generating intuitive names and adhering strictly to the predefined “Action-Objective” format.
+Method NMI (Mean $\pm$ Std) NMI Gain #Good (Mean $\pm$ Std) #Good Gain
+Baselines
+K-Means 0.7899 $\pm$ 0.0135 - 94.8% $\pm$ 1.3% -
+GMMs 0.7903 $\pm$ 0.0140 +0.05% 91.1% $\pm$ 1.6% -3.90%
+Hierarchical 0.8001 $\pm$ 0.0128 +1.29% 94.9% $\pm$ 1.2% +0.11%
+LLM-Guided Clustering Baselines
+ClusterLLM Zhang et al. (2023) 0.7284 $\pm$ 0.0168 -7.79% 91.2% $\pm$ 1.8% -3.80%
+IDAS De Raedt et al. (2023) 0.8109 $\pm$ 0.0105 +2.66% * 93.6% $\pm$ 1.2% -1.27%
+Keyphrase Viswanathan et al. (2024) 0.8371 $\pm$ 0.0098 +5.97% *** 94.5% $\pm$ 1.0% -0.32%
+LLMEdgeRefine Feng et al. (2024) 0.7411 $\pm$ 0.0155 -6.19% 87.2% $\pm$ 2.0% -8.02%
+Proposed Method (Context-Free)
+LLM-ITL (random) 0.8202 $\pm$ 0.0095 +3.84% ** 97.7% $\pm$ 0.6% +3.06% ***
+LLM-ITL (convex) 0.8208 $\pm$ 0.0090 +3.92% ** 97.8% $\pm$ 0.5% +3.16% ***
+LLM-ITL + keyphrase 0.8378 $\pm$ 0.0085 +6.06% *** 96.4% $\pm$ 0.6% +1.69% *
+Proposed Method (Context-Aware)
+LLM-ITL + merge 0.8420 $\pm$ 0.0178 +6.59% ** 97.8% $\pm$ 1.3% +3.16% ***
+LLM-ITL + role 0.8679 $\pm$ 0.0068 +9.86% *** 97.2% $\pm$ 0.5% +2.53% **
+LLM-ITL + role + merge 0.8826 $\pm$ 0.0060 +11.76% *** 97.6% $\pm$ 0.4% +2.95% ***
+Table 4: Performance of baselines and proposed methods in dialogue intent clustering. Gains are computed relative to the K-Means baseline. The best results in each column are bolded.
 5.2 Main Results
-Table 2 presents the main results evaluated using the proposed dataset.
+Table 4 reports the main results on the proposed dataset, with all evaluation metrics averaged over five random seeds.
 Among the LLM-guided clustering baselines, the data-centric keyphrase expansion method, which refines sentence embeddings via LLM-summarized keyphrases, achieved the most significant performance gain Viswanathan et al.
@@
 Interestingly, goodness evaluation does not always align with the NMI score.
-For example, while keyphrase expansion improved NMI, the generated keyphrases could distort the original sentence meaning, leading to less representative clusters and a slightly lower goodness score compared to the base model.
+For example, while keyphrase expansion improved NMI, the generated keyphrases could distort the original sentence’s meaning, leading to less representative clusters and a slightly lower goodness score compared to the base model.
 This highlights the need to balance ground truth alignment with semantic coherence during evaluation.
@@
 This emphasizes the importance of task-centric design in leveraging intent labels effectively for the problem of intent clustering.
-Notably, the proposed method does not require prior knowledge of the number of clusters and performs parameter searches during clustering, unlike baseline methods that rely on the true number of clusters.
-This adaptability underscores its strong application potential for extracting intent clusters from noisy text corpora Akama et al.
+Notably, the proposed method does not require prior knowledge of the number of clusters and automatically performs parameter searches during clustering, unlike baselines that depend on the true cluster count.
+This adaptability highlights its strong application potential for extracting intent clusters from noisy text corpora Akama et al.
 (2020).
@@
 ours 82.32 94.12 72.45 78.12
-Table 5: Performance (NMI) on English benchmarks.
+Table 5: NMI (%) performance of different clustering methods on four English intent benchmarks.
 5.3 Evaluation on English Benchmarks
@@
 Furthermore, the computational cost is lower than the compared methods as measured in Section 5.5.
+Method Accuracy Performance Gain
+K-Means (Baseline) 0.65 -
+ClusterLLM 0.62 -4.62%
+LLMEdgeRefine 0.63 -3.08%
+IDAS 0.68 +4.62% *
+Keyphrase Expansion 0.72 +10.77% ***
+LLM-ITL (context-free) 0.73 +12.31% ***
+LLM-ITL (context-aware) 0.77 +18.46% ***
+Table 6: Accuracy of BERT classifiers trained on datasets generated by different clustering methods.
 5.4 Application Performance
-One objective of intent clustering is to create a high-quality labeled dataset for training intent classifiers that can effectively handle future data inputs Gung et al.
+One objective of intent clustering is to build a high-quality labeled dataset for training intent classifiers that can handle future inputs Gung et al.
 (2023).
-To demonstrate the practical advantages of the proposed methods, we further evaluate their performance by training BERT classifiers (bert-base-chinese model) based on the clustered results.
-Compared to the training data produced by the base model, which achieved only 65% classification accuracy, the proposed LLM-in-the-loop method improved the accuracy by 12%, providing promising evidence of the improved cluster quality.
+To evaluate the practical effectiveness of our methods, we trained BERT classifiers (bert-base-chinese) on clustered data generated by different approaches.
+As shown in Table 6, the baseline method only achieved 65% accuracy.
+ClusterLLM and LLMEdgeRefine performed slightly worse, likely due to their heavier reliance on LLM-driven modifications, which can introduce label noise or overly fine-grained distinctions that reduce cluster consistency.
+In contrast, IDAS applies a more conservative refinement strategy, leading to a modest but stable improvement.
+Our proposed LLM-in-the-loop methods substantially outperformed existing approaches: LLM-ITL reached 73% accuracy, and its enhanced variant achieved 77%, representing an 18.46% relative improvement.
+These results are consistent with the observed cluster quality and provide additional empirical evidence that our method produces high-quality, human-aligned intent clusters with significant practical advantages.
 5.5 Analysis of Computational Cost
-We systematically compare the cost of the proposed method with existing LLM-guided clustering on the Bank77 dataset, with a test set of $S=3,080$ sentences and true cluster number $N=77$.
-The proposed LLM-in-the-loop iterative method (Algorithm 1) uses candidate cluster numbers $N=[10,30,50,70]$, requiring $\sum_{n_{i}\in N}n_{i}=160$ LLM calls per iteration, and with $T=3$ iterations, totals $T\cdot\sum_{n_{i}\in N}n_{i}=480$ calls, each with 20 sentences input for coherence evaluation.
-ClusterLLM, employing a fixed triplet sampling approach with $Q=1,024$, incurs a constant $1,024$ LLM calls.
-Keyphrase expansion generates one keyphrase per sentence, resulting in a linear cost of $3,080$ LLM calls for the entire dataset.
-While our method requires more input tokens per LLM call, the cost remains efficient compared to its counterparts since the output token (i.e., “good/bad” label and intent label) is kept concise with fewer LLM calls.
+We compare the computational cost of the proposed method with existing LLM-guided clustering on the Bank77 dataset, which contains $S=3,080$ test sentences and $N=77$ true clusters.
+Our LLM-in-the-loop approach evaluates candidate cluster numbers $N=[10,30,50,70]$, requiring $\sum_{n_{i}\in N}n_{i}=160$ calls per iteration.
+With $T=3$ iterations, this totals 480 calls, each processing 20 sentences for coherence evaluation.
+Cluster naming requires additional LLM calls based on the final cluster count, resulting in approximately 560 calls.
+In contrast, ClusterLLM uses a fixed triplet sampling strategy with $Q=1,024$, resulting in a constant cost of 1,024 calls.
+Keyphrase expansion generates one keyphrase per sentence, totaling 3,080 calls.
+While our method processes more input tokens per call, it remains cost-efficient as the output is limited to a short “good” or “bad” label or an intent label.
 5.6 Ablation Studies
-Effective Coherence Evaluation with Data Sampling and LLM Crowdsourcing
-Optimizing the input to LLM evaluator is crucial for enhancing performance and reducing costs Song et al.
+Data Sampling and LLM Crowdsourcing for Effective Coherence Evaluation
+Optimizing input to the LLM evaluator is crucial for improving performance and reducing costs Song et al.
 (2025); Jiang et al.
 (2024).
-As shown in Figure 3, convex sampling outperforms random sampling in selecting representative sentences for coherence evaluation.
-Interestingly, as the sample time $t$ for repeated evaluation with majority voting or the dimension of the convex hull $d$ increases, performance consistently declines.
-This suggests that simpler, more efficient parameters lead to better accuracy.
-This pattern is verified through experiments with the ChatGLM model (see red line), which exhibits the same pattern.
+In this ablation study, we compare various sampling techniques to identify best practices for coherence evaluation.
+As shown in Figure 3, convex sampling outperforms random sampling in selecting representative sentences.
+Notably, increasing either the sample repetitions $t$ for repeated validation or the convex hull dimension $d$ consistently reduces performance.
+This suggests that simpler parameter choices yield higher accuracy, a trend also seen with the ChatGLM model (red line) and summarized in Table 15 with numerical comparison.
 Figure 3: Performance comparison of sampling methods and hyperparameters for effective LLM evaluation.
-To better assess the consistency and alignment between LLM judgments and human perceptions, we utilize the concept of crowdsourcing for collaborative coherence evaluation Zhang et al.
-(2025).
+To better assess the consistency and alignment between LLM judgments and human perceptions, we utilize the concept of crowdsourcing for collaborative coherence evaluation Hong et al.
+(2025b).
 The proposed LLM crowd consists of diverse entities Zhang et al.
@@
 Based on the evaluation, we argue that a single fine-tuned LLM can effectively serve as a robust evaluator, as evidenced by the comparable performance to crowdsourced accuracy (cross markers in Figure 3) when using convex sampling.
-LLM qwen7b qwen14b baichuan2-7b chatglm3-6b
-Accuracy 62.50% 65.00% 58.50% 60.00%
-Table 6: Performance of fine-tuned LLMs in evaluating semantic coherence of intent clusters using numerical 1–4 scale scoring.
-Binary Judgment vs. Numerical Scoring
-To identify the optimal evaluation strategy for our clustering framework, we compared a numerical 1–4 scoring system (1: very poor, 4: very coherent) with the proposed binary good/bad judgment.
-We annotated 1,000 clusters with input from five human experts, finalized scores through majority voting, and fine-tuned four Chinese LLMs using this dataset.
-The models were then evaluated on 200 additional clusters, assigning scores five times per cluster and consolidating results via majority voting.
-The 1–4 scale demonstrated low accuracy and inconsistency, with 19% of samples observed to have significant disagreement (e.g., three ‘1’s and two ‘4’s) in Qwen-2.5-14B, making it unreliable for assessing cluster quality or aligning with human evaluation.
-In contrast, the simpler binary good/bad judgment provided consistent and interpretable quality assessments, proving to be the more effective evaluation strategy.
 Method NMI Goodness
@@
 Effectiveness of Cluster Merging Techniques
-Table 7 evaluates the impact of distance measures and merging decisions on the proposed cluster merging technique.
-Results highlight the superiority of geodesic distance in hyperspherical space, demonstrating its ability to effectively capture deeper semantic relationships compared to cosine similarity.
-Additionally, the probabilistic merging strategy consistently improves robustness by accounting for uncertainty, yielding better clustering quality regardless of the distance measure used.
-These findings emphasize the importance of both an appropriate distance metric and a probabilistic approach in achieving optimal merging outcomes.
-{CJK}
-UTF8gkai
-Top1 Top2 Top3 Top4 Top5
+Table 7 evaluates the effects of distance measures and merging strategies on the proposed cluster merging method.
+The results show that geodesic distance in hyperspherical space outperforms cosine similarity by capturing deeper semantic relationships.
+Furthermore, the probabilistic merging strategy consistently enhances robustness by accounting for uncertainty, leading to better clustering quality with any distance measure.
+These findings underscore the value of combining an appropriate distance metric with a probabilistic approach to achieve optimal merging performance.
+Top1
+Top2
+Top3
+Top4
+Top5
 Accuracy 26.32% 47.37% 73.68% 84.21% 89.47%
@@
 In practice, the search space for the number of clusters can be extensive, ranging from a few options to several hundred.
-To mitigate the computational costs associated with traditional model-based selection, we propose an LLM-native method for search space pruning, which predicts the optimal cluster number for subsequent iterations using logs from previous iterations (see Appendix 13).
+To mitigate the computational costs associated with traditional model-based selection, we propose an LLM-native method for search space pruning, which predicts the optimal cluster number for subsequent iterations using logs from previous iterations (see Table 16).
 By pruning the search space and selecting the top 5 non-repetitive predicted solutions, we achieved an accuracy of 89.47% (see Table 8), indicating a high likelihood that the optimal cluster number is among predicted candidates.
-This approach substantially reduces redundant model fittings and enhances the efficiency of deploying the proposed methods.
+This approximation approach significantly reduces redundant model fittings and improves the efficiency of applying the proposed methods to large-scale datasets.
+Binary Judgment vs. Numerical Scoring for Coherence Annotation
+Finally, to identify the optimal annotation strategy for coherence evaluation, we compared a numerical 1 – 4 scoring system (1: very poor, 4: very coherent) with the proposed binary good/bad judgment.
+We annotated 1,000 clusters with input from five human experts, finalized scores through majority voting, and fine-tuned four Chinese LLMs using this dataset.
+The models were then evaluated on 200 additional clusters, assigning scores five times per cluster and consolidating results via majority voting (see results in Table 9).
+The 1 – 4 scale demonstrated low accuracy and inconsistency, with 19% of samples observed to have significant disagreement (e.g., three ‘1’s and two ‘4’s) in Qwen-2.5-14B, making it unreliable for assessing cluster quality or aligning with human judgment.
+In contrast, the simpler binary good/bad judgment provided more consistent, interpretable, and reliable quality assessments, demonstrating its superiority as an evaluation protocol.
+LLM
+qwen7b
+qwen14b
+baichuan2-7b
+chatglm3-6b
+Accuracy 62.50% 65.00% 58.50% 60.00%
+Table 9: Performance of fine-tuned LLMs in assessing cluster coherence using a numerical 1 – 4 scale.
+To further validate the effectiveness and robustness of our annotation protocol, we conducted an inter-annotator agreement study with five experts.
+They annotated 100 intent clusters generated by the K-means algorithm, covering 2,000 sentences from recent customer service dialogues.
+Each cluster was evaluated using (i) a binary Good/Bad judgment and (ii) a finer-grained 1 – 4 rubric.
+The Fleiss’ kappa ($\kappa$) for the binary scheme reached 0.82, indicating “almost perfect” agreement Landis and Koch (1977), while the 1 – 4 rubric achieved only 0.59 (“moderate” agreement) due to variability in intermediate scores (2 and 3).
+These results show that binary annotation provides a more reliable and interpretable measure of cluster coherence, reinforcing the robustness of our methodology.
 6 Conclusion
-This paper tackles dialogue intent clustering using a human-aligned LLM-in-the-loop approach.
-Experiments on a complex, realistic Chinese customer service dataset reveal that fine-tuned LLM utilities excel in semantic coherence evaluation and cluster labeling, with the proposed methods consistently outperforming existing LLM-guided clustering baselines in cluster quality and computational efficiency.
-Future work should focus on refining evaluation strategies to enhance semantic guidance from LLMs and exploring task-centric approaches for deeper LLM-in-the-loop integration.
+This paper tackles dialogue intent clustering through a human-aligned LLM-in-the-loop framework.
+Experiments on a large-scale Chinese customer service dataset demonstrate that fine-tuned LLM utilities are highly effective for semantic coherence evaluation and cluster labeling, enabling consistent improvements over existing LLM-guided baselines in both clustering quality and computational efficiency.
+Beyond achieving state-of-the-art performance, our study offers strong empirical evidence for the effectiveness of LLM-in-the-loop methodologies, with ablation studies highlighting best practices.
+Future work should refine evaluation beyond coherence to capture interpretability and expressiveness, extend the clustering framework to multilingual settings, and explore deeper task-centric integration of LLMs to further advance intent mining in real-world applications.
 Limitations
-This work has two main limitations.
-First, while the coherence evaluation is practical and intuitive, relying solely on cluster coherence as the quality indicator overlooks other critical attributes, such as meaningfulness and expressiveness, which are equally important in assessing a cluster’s usefulness.
-For instance, a cluster labeled “express-feeling” may be overly general and should be refined into more specific categories like “express-appreciation” and “state-dissatisfaction” to enhance interpretability and practical applicability.
-Additionally, the binary evaluation and numerical scoring considered in this paper are deterministic indicators, overlooking probabilistic judgments and confidence intervals, which could provide deeper insights into clustering assessment and offer greater flexibility for integration into the LLM-in-the-loop clustering process.
-Second, the dataset proposed in this study is limited to the Chinese language and focuses on specific domains, namely banking, insurance, and telecommunication.
-While it represents the largest realistic customer service call dataset for the intent clustering problem, its coverage is restricted in terms of multi-domain and multilingual aspects, potentially affecting the generalizability of our findings to other languages and domains.
-Furthermore, the scope of this paper is confined to intent clustering, which serves as the initial step in the broader process of intent discovery.
-Subsequent steps, such as analyzing intent trajectories and recognizing intents at the document level rather than the sentence level, are not comprehensively discussed.
-These areas warrant further investigation to fully reveal the practical values of LLM-in-the-loop design in intent mining tasks.
+Despite the effectiveness of the proposed LLM-in-the-loop intent clustering method, this study has several limitations.
+First, while cluster coherence is a practical and intuitive quality indicator, its sole reliance overlooks other critical attributes, such as meaningfulness and expressiveness, which are equally important for assessing the quality of intent clusters.
+For example, a cluster labeled “express-feeling” may be too broad and could be refined into more specific clusters like “express-appreciation” to improve interpretability and applicability.
+Additionally, the binary evaluation and numerical scoring used are both deterministic metrics with fixed scales, neglecting probabilistic judgments and confidence intervals that could provide deeper assessment insights and enhance flexibility for LLM-in-the-loop integration.
+Second, the proposed intent clustering dataset is limited to the Chinese language and specific domains.
+Although it excels in capturing large-volume, realistic customer service dialogues, its scope restricts multi-domain and multilingual generalizability, potentially limiting the applicability of findings to other languages and domains.
+Moreover, this paper focuses solely on intent clustering, an initial step in the broader intent discovery process.
+Subsequent steps, such as analyzing intent trajectories and recognizing intents at the document level rather than the sentence level, encourage further investigation with more diverse LLM utilities and varied integration strategies to better reveal the practical value of LLM-in-the-loop methodologies.
+Acknowledgments
+This paper is partially supported by several ongoing projects led or coordinated by Prof.
+Zhang Chen, including P0045948 and P0046453 (industry donations from Accel Group Holding Limited and Minshang Creative Technology Holdings Limited), P0046701 and P0046703 (PolyU internal research funding), P0048183 and P0048191 (Research Matching Grant Scheme funded by the University Grants Committee), P0048887 (Innovation and Technology Fund - ITSP, ITS/028/22FP), P0051906 (RGC Early Career Scheme, 25600624), and P0054482 (industry donations from Two Square Capital Limited).
 Appendix A Data Collection and Annotation
-This section details the three-stage data collection and annotation process for the proposed Chinese dialogue intent clustering dataset.
+This section details the three-stage data collection and annotation process for the proposed Chinese customer service dialogue intent clustering dataset.
 Data Collection.
-The original dialogue dataset is sourced from audio transcriptions of customer service calls across several domains.
-Initially, the dataset contains over 11,879 customer service calls.
-Strict filtering is applied to maintain data integrity and exclude confidential or sensitive content, resulting in 8,184 dialogues and 69,839 sentences.
-Subsequent data cleaning removes duplicate sentences, resulting in a final set of 55,085 unique sentences.
+The raw dialogue dataset was derived from audio transcriptions of customer service calls in three major domains: banking, telecommunication, and insurance.
+In total, the dataset contained 11,879 calls.
+To ensure data integrity and protect confidentiality, strict filtering was applied to exclude sensitive content, resulting in 8,184 dialogues with 69,839 sentences.
+Further cleaning was conducted to remove duplicate sentences, yielding a final set of 55,085 unique sentences.
 Data Annotation.
-To construct the intent clustering dataset, we internally recruited 15 human experts, each with over five years of experience in customer service call centers, to participate in the data annotation process:
+To construct the intent clustering dataset, we recruited 15 human experts to conduct the annotations.
+Each annotator has over five years of professional experience in customer service call centers, with domain expertise aligned with the dataset domains: banking, telecommunication, and insurance.
+All annotators are fluent in Chinese, familiar with domain-specific terminology, and experienced in dialogue annotation or quality assurance tasks.
+Prior to annotation, a mandatory training session was held, during which the authors provided 50 intent clusters annotated with Good/Bad judgments and Action–Objective intent labels as demonstration and reference examples.
+This training ensured a consistent understanding of the guidelines across all annotators.
+The annotation process is outlined as follows:
   1. 1.
@@
   3. 3.
-    For the good clusters, annotators were asked to label the underlying intent using the naming convention “Action-Objective.” Five samples were provided as guides.
-After intent annotation, 1255 clusters with unique intentions were retained, and 28 clusters were merged due to replicated intent.
+    For the good clusters, annotators were asked to label the underlying intent using the naming convention “Action-Objective.” After intent annotation, 1255 clusters with unique intentions were retained, and 28 clusters were merged due to replicated intent.
   4. 4.
@@
 Example: input:[sentences] output:[label] input:{[sentences]} output:
+Model Coherence Evaluation Cluster Naming
+LLaMA-2-7B (LoRA) 94.0% $\pm$ 1.2% 89.0% $\pm$ 1.5%
+LLaMA-2-13B (LoRA) 95.0% $\pm$ 1.1% 93.5% $\pm$ 1.4%
+Mistral-7B (LoRA) 91.5% $\pm$ 2.3% 82.0% $\pm$ 1.9%
+GPT-3.5 (API) 91.0% $\pm$ 1.5% 86.0% $\pm$ 1.8%
+GPT-4 (API) 94.5% $\pm$ 1.2% 91.0% $\pm$ 1.3%
+Table 10: Performance of different LLMs on coherence evaluation and cluster naming for the English dataset.
 Appendix C More Results: Evaluation with Proprietary LLM
-Given the resource-intensive nature of LLM fine-tuning, we conducted further experiments in coherence evaluation using proprietary LLMs accessed through OpenAI APIs: GPT-3.5, GPT-4, and GPT-4o.
-These advanced and widely used models offer strong performance in various tasks but also come with higher operational costs due to token consumption.
-Using the same evaluation set of 480 clusters, these models achieved accuracies of 89.58%, 93.54%, and 94.17%, respectively.
-The results highlight that a smaller, fine-tuned LLM, such as Qwen-2.5-7b (with an accuracy of 96.25%), can outperform these advanced models while significantly reducing costs associated with API usage [^4] The fine-tuned Chinese LLMs for coherence evaluation and cluster naming will be released upon acceptance..
-For comparison, a vanilla Qwen-2.5-7b model (without fine-tuning) was also evaluated, yielding a substantially lower accuracy of 75.63%, emphasizing the essence of fine-tuning.
-{CJK}
-UTF8gkai
-Cluster Coherence Original Sentences English Translation
-Good "给企业固定资产买保险,大约能投大约得投保多少呢", "就是假如我有一百万的企业固定资金买保险大概投保多少钱啊", "请问一下我想咨>询一下企业财产保险", "您好我想了解一下这个企业财产保险", "就了解一下这个企业财产保险","你们是财险", "是哦这个属于财险了对吧", "企业财产保险的", "这些都属于财产险对吗", "财险人工那你这是", 就是如果我要为我这个私营企业买这个保险需要什么手续", "财产险", "呃企业财产保险是以什么为保" "If I buy insurance for a company’s fixed assets, how much insurance will it cost?", "That is, if I have a million corporate fixed assets, how much will it cost to buy insurance?", "Excuse me, I would like to consult> Ask about corporate property insurance", "Hello, I want to know about this corporate property insurance", "Just want to know about this corporate property insurance", "You are a property insurance company", "Yes, this belongs to property insurance "Right?", "Enterprise property insurance", "These all belong to property insurance, right?", "What about property insurance workers?", What do I need if I want to buy this insurance for my private enterprise?
-Procedure","Property Insurance", "Well, what does corporate property insurance cover?"
-Bad "就是连连续,就是一直一直保", "但是它连不上是怎么回事啊?", "哦就是主要是直接给公司转账对吧?", "接吗", "接也是吗?", "我直>接去", "你是直接直接用那个" , "就是从哪接过来再接回去","还是需要从哪儿连线这个宽带", "你直接给我说这些啊", "直接把", "哦直接", "嗯那个礼品是直接就发放了", "是你直接给我回复对吗", "直接到那里去", "嗯,最好直飞。", "请帮我连接+", "把钱打过去的话,我是直接打到那个证券公司", "直接就是" "It’s continuous, it’s always guaranteed", "But what’s wrong with it not being able to connect?", "Oh, it’s mainly to transfer money directly to the company, right?", "Yes", "Yes too?" ?", "I’ll go directly", "Are you using that directly" , "Just connect it from where you are and then connect it back", "Or do you need to connect to the broadband from somewhere", "You Just tell me this directly", "Just give it directly", "Oh directly", "Well, the gift was given out directly", "You replied to me directly, right?","Go there directly" , "Well, it’s best to fly directly.", "Please help me connect +", "If I call the money, I will call the securities company directly", "Directly"
+Given the resource-intensive nature of LLM fine-tuning, we conducted additional experiments using proprietary LLMs accessed through OpenAI APIs: GPT-3.5, GPT-4, and GPT-4o.
+These widely adopted models deliver strong performance across diverse tasks and do not require task-specific fine-tuning, thereby alleviating the data scarcity issue, but they also incur higher operational costs due to token consumption.
+On the same Chinese coherence evaluation dataset with 480 clusters, these models achieved accuracies of 89.58%, 93.54%, and 94.17%, respectively.
+Notably, the smaller fine-tuned Qwen-2.5-7B reached 96.25% accuracy, surpassing these advanced proprietary models while significantly reducing API-related costs.
+For comparison, the vanilla Qwen-2.5-7B model (without fine-tuning) obtained a much lower accuracy of 75.63%, further underscoring the importance of fine-tuning.
+On the English dataset, we report results for GPT-3.5 and GPT-4, as well as two additional fine-tuned models: LLaMA-2-13B (LoRA) and Mistral-7B (LoRA).
+The LLaMA and Mistral models were fine-tuned on the same 800 intent clusters as LLaMA-2-7B in Section 5.3 and evaluated on 100 English intent clusters annotated by human experts with Good/Bad and Action-Objective intent labels.
+Each model was run five times with different random seeds to ensure robust performance metrics.
+Results in Table 10 show that scaling up model size (e.g., 7B vs.
+13B) improves performance, while fine-tuned smaller LLMs often outperform large proprietary LLMs, consistent with the findings on the Chinese dataset.
+These results reinforce the value of fine-tuning and suggest that smaller, cost-efficient models can play a critical role in data mining within the LLM-in-the-loop framework.
+Cluster Coherence
+Original Sentences
+English Translation
+Good
+”给企业固定资产买保险,大约能投大约得投保多少呢”,
+”就是假如我有一百万的企业固定资金买保险大概投保多少钱啊”,
+”请问一下我想咨¿询一下企业财产保险”,
+”您好我想了解一下这个企业财产保险”,
+”就了解一下这个企业财产保险”,”你们是财险”,
+”是哦这个属于财险了对吧”, ”企业财产保险的”,
+”这些都属于财产险对吗”, ”财险人工那你这是”,
+就是如果我要为我这个私营企业买这个保险需要什么手续”,
+”财产险”, ”呃企业财产保险是以什么为保”
+”If I buy insurance for a company’s fixed assets,
+how much insurance will it cost?”,
+”That is, if I have a million corporate fixed assets,
+how much will it cost to buy insurance?”,
+”Excuse me, I would like to consult¿ Ask about corporate property insurance”,
+”Hello, I want to know about this corporate property insurance”,
+”Just want to know about this corporate property insurance”,
+”You are a property insurance company”,
+”Yes, this belongs to property insurance ”Right?”, ”Enterprise property insurance”,
+”These all belong to property insurance, right?”,
+”What about property insurance workers?”,
+What do I need if I want to buy this
+insurance for my private enterprise? Procedure”,”Property Insurance”,
+”Well, what does corporate property insurance cover?”
+Bad
+”就是连连续,就是一直一直保”, ”但是它连不上是怎么回事啊?”,
+”哦就是主要是直接给公司转账对吧?”, ”接吗”, ”接也是吗?”,
+”我直¿接去”, ”你是直接直接用那个”
+, ”就是从哪接过来再接回去”,”还是需要从哪儿连线这个宽带”,
+”你直接给我说这些啊”,
+”直接把”, ”哦直接”, ”嗯那个礼品是直接就发放了”,
+”是你直接给我回复对吗”, ”直接到那里去”,
+”嗯,最好直飞。”, ”请帮我连接+”,
+”把钱打过去的话,我是直接打到那个证券公司”, ”直接就是”
+”It’s continuous, it’s always guaranteed”,
+”But what’s wrong with it not being able to connect?”,
+”Oh, it’s mainly to transfer money directly to the company, right?”,
+”Yes”, ”Yes too?” ?”,
+”I’$l$ go directly”, ”Are you using that directly”
+, ”Just connect it from where you are and then connect it back”,
+”Or do you need to connect to the broadband from somewhere”,
+”You Just tell me this directly”, ”Just give it directly”, ”Oh directly”,
+”Well, the gift was given out directly”,
+”You replied to me directly, right?”,”Go there directly” ,
+”Well, it’s best to fly directly.”, ”Please help me connect +”,
+”If I call the money, I will call the securities company directly”, ”Directly”
 Table 11: Example of “Good” and “Bad” intent clusters in coherence evaluation.
-{CJK}
-UTF8gkai
-Cluster Name Original Sentences English Translation
-询问-优惠 (Inquire-Promotion) ’那有什么优惠券什么之类的吗?’, ’是是怎么形式是优惠吗?’, ’还是不是优惠活动,是那个直接给我打我卡里吗 ,还是是什么优惠券儿啊?’, ’就比如新用户他有什么优惠券儿之类的吧!’, ’嗯你你那还有什么优惠的活动吗?
-就是比较合适就是合适的动。’ ’Are there any coupons or anything like that?
-’, ’What is the form of the discount?
-’, ’Still, it’s not a promotion.
-Is it the one that directly charges my card?
-, or is it some kind of coupon?
-’, ’For example, what coupons does a new user have?
-’, ’Well, do you have any other discounts?
-It is more appropriate and appropriate to move.
-’
-解答-金额 (Answer-Amount) ’一共是三十一块二毛’, ’就每个月一百三十八’, ’对,一个月也就是四百百四五百块钱嘛, 给您自己做个积累。’, ’十二月份的话是用了三十三块九毛二’, ’对一个月一百三十八’ ’The total is thirty-one and twenty cents’, ’That’s one hundred and thirty-eight cents per month’, ’Yes, that’s four hundred, four hundred and five hundred yuan per month.
-Make an accumulation for yourself.
-’, ’In December, it cost thirty-three dollars and ninety-two cents’, ’That’s one hundred and thirty-eight dollars a month’
-Table 10: Example of cluster naming with "Action-Objective" style.
-Dataset Naming Convention Example (sentence) Example (label)
-NLU scenario-intent Send an email to Alex and write thank you. email sendemail
-NLU++ list of keywords How long does it usually take to get a new pin? ["how_long","pin", "arrival","new"]
-OOS objective Please tell me why my card was declined yesterday. card_declined
-ours action-objective Well, do you have any other discounts? inquire-promotion
-Table 11: Comparison of cluster naming convention in intent clustering datasets.
-{CJK}
-UTF8gkai
-Cluster Name Clusters with Similar Intention English Translation
-询问-意外事故 (Inquire-Accident) "假如被车碰了或者是被楼上的砖砸了一下", "给别人儿撞的意外", "啊撞到别人然后就是", "就平时有时候开车嘛可能会遇到这个", "嗯哦这种情况,那要是就是我自己不小心撞到了那个某个地方然后", "就是把其他的东西撞到了呀什么的", "那人生意,不是就是,如果是不小心在马上被车撞了的话", "撞到人了吧", "然后不小在行驶当中被别人损害就是说拿石头砸的呀然后", "什么被车撞了之类的,是吗", "被被被撞了,还是被什么一些什么意外事故了", "我把别人的车撞了是吧" "If I were hit by a car or hit by a brick from upstairs," "Accidentally hit by someone else," "Ah, hit someone else and then," "Just sometimes when driving, you might encounter this," "Well, if I’m not careful and hit some place myself," "Just hit something else or something," "That’s a human affair, not just, if it’s accidentally hit by a car on the road," "Hit someone, right?" "Then not small in the process of driving, being damaged by someone else, say, hit with a stone, and then," "What, hit by a car or something, right?" "I hit someone else’s car, right?"
-询问-意外死亡 (Inquire-Accident Death) "哦猝死,那猝死算意外吗", "啊,那我想知道那个猝死的话,算是意外死亡吗", "算意外死亡吗", "那如果猝死算是意外死亡吗?", "猝死也算意外事吧", "那那猝死是意外死亡吗如果是猝死的话是", "猝死算是意外死亡吗", "猝死算意外死亡吗", "嗯,那个猝死,猝死属于意外意外险吗,意外死亡吗", "那猝死的话,算意外死亡吗?", "那那那那个就是那个猝死算是意外死亡吗" "Oh sudden death, is sudden death considered accidental?" "Ah, I want to know if sudden death is considered accidental death?" "Is it considered accidental death?" "If sudden death is considered accidental death?" "Sudden death is also an accident, right?" "Is sudden death considered accidental death if it is sudden death?" "Is sudden death considered accidental death?" "Is sudden death considered accidental death?" "Well, that sudden death, does sudden death fall under accidental insurance, accidental death?" "Is sudden death considered accidental death?" "Is that, that, that sudden death considered accidental death?"
+Cluster Name
+Original Sentences
+English Translation
+询问-优惠
+(Inquire-Promotion)
+’那有什么优惠券什么之类的吗?’,
+’是是怎么形式是优惠吗?’,
+’还是不是优惠活动,是那个直接给我打我卡里吗
+,还是是什么优惠券儿啊?’,
+’就比如新用户他有什么优惠券儿之类的吧!’,
+’嗯你你那还有什么优惠的活动吗?
+就是比较合适就是合适的动。’
+’Are there any coupons or anything like that? ’,
+’What is the form of the discount? ’,
+’Still, it’s not a promotion. Is it the one that directly charges my card?
+, or is it some kind of coupon? ’,
+’For example, what coupons does a new user have? ’,
+’Well, do you have any other discounts?
+It is more appropriate and appropriate to move. ’
+解答-金额
+(Answer-Amount)
+’一共是三十一块二毛’,
+’就每个月一百三十八’,
+’对,一个月也就是四百百四五百块钱嘛,
+给您自己做个积累。’,
+’十二月份的话是用了三十三块九毛二’,
+’对一个月一百三十八’
+’The total is thirty-one and twenty cents’,
+’That’s one hundred and thirty-eight cents per month’,
+’Yes, that’s four hundred,
+four hundred and five hundred yuan per month. Make an accumulation for yourself. ’,
+’In December, it cost thirty-three dollars and ninety-two cents’,
+’That’s one hundred and thirty-eight dollars a month’
+Table 12: Example of cluster naming with ”Action-Objective” convention.
+Dataset
+Naming Convention
+Example (sentence)
+Example (label)
+NLU
+scenario-intent
+Send an email to Alex and write thank you.
+email sendemail
+NLU++
+list of keywords
+How long does it usually take to get a new pin?
+[”how_long”,”pin”,
+”arrival”,”new”]
+OOS
+objective
+Please tell me why my card was declined yesterday.
+card_declined
+ours
+action-objective
+Well, do you have any other discounts?
+inquire-promotion
+Table 13: Comparison of cluster naming conventions in existing and proposed intent clustering datasets.
+Cluster Name
+Clusters with Similar Intention
+English Translation
+询问-意外事故
+(Inquire-Accident)
+”假如被车碰了或者是被楼上的砖砸了一下”,
+”给别人儿撞的意外”,
+”啊撞到别人然后就是”,
+”就平时有时候开车嘛可能会遇到这个”,
+”嗯哦这种情况,那要是就是我自己不小心撞到了那个某个地方然后”,
+”就是把其他的东西撞到了呀什么的”,
+”那人生意,不是就是,如果是不小心在马上被车撞了的话”,
+”撞到人了吧”,
+”然后不小在行驶当中被别人损害就是说拿石头砸的呀然后”,
+”什么被车撞了之类的,是吗”,
+”被被被撞了,还是被什么一些什么意外事故了”,
+”我把别人的车撞了是吧”
+”If I were hit by a car or hit by a brick from upstairs,”
+”Accidentally hit by someone else,”
+”Ah, hit someone else and then,”
+”Just sometimes when driving, you might encounter this,”
+”Well, if I’m not careful and hit some place myself,”
+”Just hit something else or something,”
+”That’s a human affair, not just, if it’s accidentally
+hit by a car on the road,”
+”Hit someone, right?”
+”Then not small in the process of driving, being
+damaged by someone else, say, hit with a stone, and then,”
+”What, hit by a car or something, right?”
+”I hit someone else’s car, right?”
+询问-意外死亡
+(Inquire-Accident Death)
+”哦猝死,那猝死算意外吗”,
+”啊,那我想知道那个猝死的话,算是意外死亡吗”,
+”算意外死亡吗”,
+”那如果猝死算是意外死亡吗?”,
+”猝死也算意外事吧”,
+”那那猝死是意外死亡吗如果是猝死的话是”,
+”猝死算是意外死亡吗”,
+”猝死算意外死亡吗”,
+”嗯,那个猝死,猝死属于意外意外险吗,意外死亡吗”,
+”那猝死的话,算意外死亡吗?”,
+”那那那那个就是那个猝死算是意外死亡吗”
+”Oh sudden death, is sudden death considered accidental?”
+”Ah, I want to know if sudden death is considered accidental death?”
+”Is it considered accidental death?”
+”If sudden death is considered accidental death?”
+”Sudden death is also an accident, right?”
+”Is sudden death considered accidental death if it is sudden death?”
+”Is sudden death considered accidental death?”
+”Is sudden death considered accidental death?”
+”Well, that sudden death, does sudden death fall under
+accidental insurance, accidental death?”
+”Is sudden death considered accidental death?”
+”Is that, that, that sudden death considered accidental death?”
 Table 14: Example of two high-quality intent clusters with similar intentions.
+Model Sampling Method 1 2 3 4 5
+qwen14b convex - 96.25% 95.42% 95.42% 95.21%
+random (n=10) 95.83% 94.38% 93.33% 92.92% 92.92%
+random (n=20) 94.58% 95.00% 94.38% 94.17% 94.17%
+chatglm3-6b convex - 92.08% 90.83% 91.04% 90.42%
+random (n=10) 92.29% 88.75% 85.83% 84.58% 85.21%
+random (n=20) 90.42% 89.58% 89.38% 89.58% 88.98%
+Table 15: Comparison of sampling methods and hyperparameters for LLM coherence evaluation.
 Epoch 1th 2th 3th
@@
 Best 1600 52184 2901 17.988 800 2454 447 5.490 200 86 361 0.238
-Table 13: Example of log record for iterative intent clustering.
+Table 16: Example log records from iterative intent clustering across epochs.